### v1

# TSPRank: Bridging Pairwise and Listwise Methods with a Bilinear Travelling Salesman Model

Abstract.
Traditional Learning-To-Rank (LETOR) approaches, including pairwise methods like RankNet and LambdaMART, often fall short by solely focusing on pairwise comparisons, leading to sub-optimal global rankings. Conversely, deep learning based listwise methods, while aiming to optimise entire lists, require complex tuning and yield only marginal improvements over robust pairwise models. To overcome these limitations, we introduce Travelling Salesman Problem Rank (TSPRank), a hybrid pairwise-listwise ranking method. TSPRank reframes the ranking problem as a Travelling Salesman Problem (TSP), a well-known combinatorial optimisation challenge that has been extensively studied for its numerous solution algorithms and applications. This approach enables the modelling of pairwise relationships and leverages combinatorial optimisation to determine the listwise ranking. This approach can be directly integrated as an additional component into embeddings generated by existing backbone models to enhance ranking performance. Our extensive experiments across three backbone models on diverse tasks, including stock ranking, information retrieval, and historical events ordering, demonstrate that TSPRank significantly outperforms both pure pairwise and listwise methods. Our qualitative analysis reveals that TSPRank’s main advantage over existing methods is its ability to harness global information better while ranking. TSPRank’s robustness and superior performance across different domains highlight its potential as a versatile and effective LETOR solution. The code and preprocessed data are available at https://github.com/waylonli/TSPRank-KDD2025.
learning-to-rank, pairwise-listwise ranking, travelling salesman problem
copyright: acmlicensedconference: Proceedings of the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, Canadabooktitle: Proceedings of the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’25), August 3–7, 2025, Toronto, Canadajournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Learning to rank
1. Introduction
Learning to Rank (LETOR) algorithms have become essential in applications such as recommendation systems (Fan et al., 2022; Pei et al., 2019; Li et al., 2021), question answering (Deng et al., 2021; Ji and Wang, 2013; Campese et al., 2024), and information retrieval (Ai et al., 2018; Pang et al., 2020; Buyl et al., 2023). These algorithms aim to order a list of ranking entities based on their features, optimising for the most relevant or preferred entities to appear at the top. Recently, LETOR methods have also expanded into other domains, such as stock and portfolio selection (Song et al., 2017; Wu et al., 2024) and textual ordering (Tan et al., 2013; Wang et al., 2023). Despite these broader applications of LETOR, latest fundamental work on LETOR still primarily focused on incorporating click data into ranking models and addressing biases introduced by user feedback (Oosterhuis, 2023; Gupta et al., 2024; Zhang et al., 2023; Guo et al., 2020), concentrating on retrieval and recommendation tasks. Limited research has explored general ranking methods across diverse tasks and domains.
Over the past decade, research on LETOR models mainly focuses on pairwise and listwise approaches, while pointwise methods often fail to capture the intricate inter-entity relationships that are essential for accurate ranking. Pairwise methods, such as LambdaMART and RankNet (Burges, 2010), primarily optimise for pairwise comparisons without a holistic view of the entire ranking list, potentially leading to sub-optimal global rankings. Listwise methods optimise the ranking of entire lists directly rather than individuals or pairs. This category includes advanced neural network architectures, particularly adaptations of the Transformer architecture (Vaswani et al., 2017), such as Rankformer (Buyl et al., 2023) and SetRank (Pang et al., 2020). However, existing work shows that deep learning-based listwise models require complex tuning to achieve marginal gains over robust pairwise models like LambdaMART on information retrieval benchmarks (Qin et al., 2021). Therefore, pairwise and listwise methods each have inherent drawbacks and more effective solutions for LETOR have not been exhaustively explored.
Recently, GNNRank successfully used directed graph neural networks to learn listwise rankings from pairwise comparisons (He et al., 2022). Although their method is not applicable to general LETOR problems, as it requires pre-known pairwise relationships like the outcomes of sports matches, it still suggests a potential approach for LETOR by representing pairwise comparisons as a graph. Inspired by Abend et al., who used combinatorial optimisation techniques to recover the lexical order of recipe pieces (Abend et al., 2015), we propose Travelling Salesman Problem Rank (TSPRank). This hybrid pairwise-listwise method reframes the ranking problem as a Travelling Salesman Problem (TSP), an NP-hard combinatorial optimisation challenge that seeks the optimal sequence of node visits to minimise total travel cost. In a ranking context, TSPRank aims to find the optimal permutation of ranking entities to maximise the ranking score.
Our main contribution, TSPRank, bridges the gap between pairwise and listwise ranking models through combinatorial optimisation, demonstrating strong robustness across diverse ranking problems. TSPRank simplifies the complex listwise ranking problem into easier pairwise comparisons, thus overcoming the challenges of directly learning listwise rankings in complex data and tasks. It also enhances pairwise methods by incorporating listwise optimisation through a TSP solver. To improve ranking performance, TSPRank can be applied directly as an additional component on embeddings generated from domain-specific backbone models. To our knowledge, it is the first LETOR model to frame the ranking problem as a combinatorial optimisation plus graph representation task. We demonstrate its superior performance across three backbone models on diverse ranking tasks: stock ranking, information retrieval, and historical events ordering, covering both numerical and textual data across multiple domains. We introduce two learning methods for TSPRank: a local method using a pre-defined ground truth adjacency matrix without the TSP solver during training and a global end-to-end method that includes the solver in the training loop for better model-inference alignment. Our empirical analysis visualises the predictions and graph connections to provide insights into TSPRank’s effectiveness. We demonstrate that the TSP solver enhances the model’s ability to harness global information during ranking and increases tolerance to errors in pairwise comparisons. Additionally, we assess inference latency caused by the combinatorial optimisation solver and suggest potential solutions to mitigate this overhead.
2. Background
LETOR algorithms aim to rank entities based on their features. In LETOR, we denote a list of $N$ ranking entities as $\mathbf{e}$, containing $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{N}\}$, with each $\mathbf{e}_{i}\in\mathbb{R}^{d}$ where $d$ represents the dimensionality of the feature space. A scoring function $s:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is applied on the entities, which are then ranked in descending order of the $s(\mathbf{e}_{i})$ scores. LETOR algorithms are broadly categorised into pointwise, pairwise, and listwise approaches based on their optimisation strategies. Pointwise methods treat ranking as a regression problem, comparing scores $s(\mathbf{e}_{i})$ to labels $y_{i}$. Models such as OPRF (Fuhr, 1989), TreeBoost (Friedman, 2001), and RankSVM (Shashua and Levin, 2002), including newer implementations like RankCNN (Severyn and Moschitti, 2015), are known for their computational efficiency but may not capture inter-entity relationships adequately.
Pairwise methods focus on the relative comparisons of entity pairs, assessing scores $(s(\mathbf{e}_{i}),s(\mathbf{e}_{j}))$ against binary labels $y_{ij}$, which denote if entity $\mathbf{e}_{i}$ should be ranked higher than $\mathbf{e}_{j}$. Methods in this category include RankNet (Burges et al., 2005), LambdaRank (Burges et al., 2006), and LambdaMART (Burges, 2010). Listwise methods directly compute scores based on the features of all entities in a list. These methods aim to optimise a listwise objective, comparing the entire set of computed scores $s(\mathbf{e})\in\mathbb{R}^{N}$ against a complete set of labels $\mathbf{y}\in\mathbb{R}^{N}$. To advance this approach, Ai et al. introduced the Deep Listwise Context Model (DLCM) (Ai et al., 2018). This model was further refined into general multivariate scoring functions in (Ai et al., 2019). Considering that a list of entities can be viewed as a sequence, methodologies from Natural Language Processing (NLP) are applicable to ranking tasks. Notably, the Transformer architecture (Vaswani et al., 2017), has been adapted to address listwise ranking challenges, referred to as listwise Transformers in recent work (Pei et al., 2019; Pasumarthi et al., 2020; Pang et al., 2020; Zhuang et al., 2021; Li et al., 2022; Buyl et al., 2023).
Although listwise ranking methods, primarily neural rankers, have demonstrated their advantages in web search reranking tasks (Ai et al., 2018; Pang et al., 2020), Qin et al. conducted a comprehensive evaluation comparing listwise neural rankers to the gradient-boosted-decision-tree (GBDT)-based pairwise LambdaMART (Burges, 2010; Qin et al., 2021). The results indicated that LambdaMART consistently outperforms listwise rankers across three widely used information retrieval datasets. Furthermore, it was observed that listwise rankers require substantial efforts to achieve only marginal improvements over LambdaMART. Consequently, it is evident that listwise ranking models are not universally applicable solutions and generally lack the robustness needed for broader deployment. This conclusion has also been further validated by Buyl et al. (Buyl et al., 2023).
3. TSPRank
Directly predicting the listwise order is challenging, as correctly ranking $N$ entities from 1 to $N$ is complex. However, breaking it down into $N\times N$ pairwise comparisons simplifies the task, as each pairwise comparison is more straightforward than the entire listwise ranking. However, a fundamental drawback of pairwise ranking, as highlighted by Cao et al. (Cao et al., 2007), is that the learning objectives do not focus on minimising the listwise ranking errors. To address this gap and rethink the approach of pairwise ranking, we introduce TSPRank, a novel methodology for position-based ranking problems. This approach integrates listwise optimisation into pairwise comparisons by modelling ranking tasks as TSP. Notably, TSPRank can be applied as an additional component with embeddings generated from domain-specific backbone models to achieve improved ranking performance.
3.1. Ranking As Travelling Salesman Problem
We model the ranking problem as an open-loop Travelling Salesman Problem (TSP). Given a set of entities ${\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{N}}$ and all the pairwise score values $s(\mathbf{e}_{i},\mathbf{e}_{j})$ indicating the gain of ranking entity $\mathbf{e}_{j}$ immediately after $\mathbf{e}_{i}$, we define a complete graph $G=(V,W)$, where $V$ represents the entities and $W$ represents the pairwise scores between every two entities. Each entity $\mathbf{e}_{i}$ in the ranking problem corresponds to a city in the TSP. By mapping entities to cities and pairwise scores to weights, we can formulate the ranking problem as a TSP to find the optimal permutation $\pi$. Thus, the objective is to find an optimal open-loop tour $\pi^{*}$ of the entities defined in Equation 1 such that the total pairwise score is maximised.
[…]
The next step is developing a scoring function that accurately models the pairwise relationships between the ranking entities, thereby effectively determining the scores.
3.2. Scoring Model
The scoring model takes a set of ranking entities or their corresponding embeddings $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{N}\}$, where each $\mathbf{e}_{i}\in\mathbb{R}^{d}$ and $d$ is the dimension of the features or embeddings. It outputs an $N\times N$ weighted adjacency matrix $A$ that represents the pairwise scores of every two entities. The scoring model comprises an optional node encoder and a trainable bilinear model on top.
Encoder
The default encoder used is the Transformer encoder block (Vaswani et al., 2017), but it can be replaced with any other encoder suited to the specific task and dataset. Additionally, using an encoder is optional if sufficiently robust backbone models are available.
Trainable Bilinear Model
Given a pair of encoded representations of entities $(\mathbf{e}_{i},\mathbf{e}_{j})$, the bilinear model computes the pairwise score using the following bilinear form:
(2) $s(\mathbf{e}_{i},\mathbf{e}_{j})=\mathbf{e}_{i}^{\intercal}\mathbf{W}\mathbf{e}_{j}+b,$
where $\mathbf{W}$ and $b$ are learnable parameters that can be optimised along with the encoder parameters.
The adjacency matrix $A$ is constructed as follows:
(3) $A_{ij}=s(\mathbf{e}_{i},\mathbf{e}_{j}),\ \text{for all }i,j\in\{1,\dots,N\},$
where each entry $A_{ij}$ directly corresponds to the computed pairwise comparison score from entity $\mathbf{e}_{i}$ to entity $\mathbf{e}_{j}$.
3.3. Inference
Inference of TSPRank, described in Equation 1, requires solving the open-loop TSP, which is widely recognised as an NP-hard combinatorial optimisation problem. The solution algorithms of the TSP consist of exact algorithms (brunch and bound (Poikonen et al., 2019)) and approximation algorithms (ant colony optimisation (Yang et al., 2008)). Given the scale of the ranking problem and the desired ranking accuracy, we prefer to obtain an exact solution for the TSP rather than an approximate one. Therefore, we formulate the TSP as a Mixed Integer Linear Programming (MILP) problem (Rieck and Zimmermann, 2010), where the integer variables represent binary decisions on whether a path between two nodes is included in the optimal tour. We then use an MILP optimization solver, such as CPLEX (Cplex, 2009) or Gurobi (Gurobi Optimization, LLC, 2023), to find the exact optimal solution.
The decision variables are defined as follows:
(4) $x_{ij}=\begin{cases}1,&\text{if entity $\mathbf{e}_{j}$ is ranked immediately after $\mathbf{e}_{i}$.}\\ 0,&\text{otherwise.}\end{cases}$
We then apply the TSP optimisation problem defined in Appendix A to ensure the result is a single, complete ranking that includes all entities. This inference process effectively performs listwise optimisation, as it determines the optimal permutation of entities that maximises the overall pairwise ranking score.
4. Learning
We introduce two distinct learning methodologies for the TSPRank model outlined in Section 3. The first method is a local approach, focusing exclusively on individual pairwise comparisons during the training phase. The second method is a global, end-to-end approach, incorporating the Gurobi solver directly into the training process, thereby allowing the black-box MILP solver to influence the learning dynamically. Figure 1 illustrates the complete pipeline of the two learning methods as well as the model architecture.
Figure 1. Illustration of TSPRank and the complete pipeline of local and global learning. The pipeline starts with a Transformer Encoder or any embeddings generated from another existing Backbone Model, followed by a Bilinear Model generating pairwise scores to form an Adjacency Matrix. Local learning compares the predicted matrix with the ground truth to calculate the local loss $\mathcal{L}_{local}$. Global learning uses the max-margin loss $\mathcal{L}_{glob}$ to incorporate the Gurobi TSP solver during training.
4.1. Local Learning
In the local learning approach, the objective is to understand pairwise consecutive relationships between nodes. The objective is to determine whether entity $\mathbf{e}_{j}$ should rank one position after $\mathbf{e}_{i}$ in a given pair of nodes. Therefore, the score $s(\mathbf{e}_{i},\mathbf{e}_{j})$ is set high if $\mathbf{e}_{j}$ is supposed to rank consecutively after $\mathbf{e}_{i}$, and low otherwise.
To tailor the TSPRank model to ranking scenarios where penalties vary based on the actual positions, we apply a weighted loss. We denote the predicted adjacency matrix $A^{p}$ from the bilinear model, where $a_{ij}^{p}\in A^{p}$ represents the predicted pairwise scores. The target adjacency matrix $A^{t}=\{a^{t}_{ij}\}$ is defined such that $a_{ij}^{t}\in A^{t}$ equals 1 if $\mathbf{e}_{j}$ ranks immediately after $\mathbf{e}_{i}$, and 0 otherwise. We apply the weighted cross-entropy loss (defined in Equation 5) on each row to identify the next entity $\mathbf{e}_{\pi(i+1)}$ to be ranked given $\mathbf{e}_{\pi(i)}$. The model is trained as a multi-class classification problem, aiming to maximise the probability of $P(\mathbf{e}_{\pi(i+1)}\mid\mathbf{e}_{\pi(i)})$. We weight the loss by the true ranking label $y$, allowing the penalty to vary according to different ranking positions. Additionally, $y$ can be adjusted to $N+1-y$ depending on whether $y$ represents ascending or descending order.
(5) $\mathcal{L}_{local}(A^{p},A^{t})=-\sum_{i=1}^{N}y_{k}\log\frac{e^{A^{p}_{ik}}}{\sum_{j=1}^{N}e^{A^{p}_{ij}}},\ k=\operatorname*{arg\,max}_{j}A^{t}_{ij}.$
4.2. Global Learning
We expect that a globally trained TSPRank model will further enhance the performance as the local learning method focuses solely on pairwise comparisons in isolation during training. Incorporating the TSP solver in the training procedure will better align the model with the inference process. Therefore, we introduce an end-to-end approach. After obtaining the predicted adjacency matrix $A^{p}$, defined in Section 4.1, we define the gold decision variable $\mathbf{x}^{t}=\{x^{t}_{ij}\}$, where $x^{t}_{ij}$ is defined as in Equation 4. We aim to train a model to satisfy the margin constraints (Equation 6) for all the possible decision variables $\mathbf{x}$ by minimising the max-margin loss defined in Equation 7. When $\mathbf{x}=\mathbf{x}_{t}$, the predicted ranking perfectly aligns with the gold ranking, resulting in a loss of zero.
(6) $\mathbf{x}^{t}\cdot A^{p}$ $\geq\mathbf{x}\cdot A^{p}+\Delta(\mathbf{x}^{t},\mathbf{x}),\text{ for all }\mathbf{x}\text{,}$
(7) $\mathcal{L}_{glob}(A^{p},\mathbf{x}^{t})$ $=\max(0,\max_{\mathbf{x}\neq\mathbf{x}^{t}}[\Delta(\mathbf{x},\mathbf{x}^{t})+\mathbf{x}\cdot A^{p}]-\mathbf{x}^{t}\cdot A^{p})$
The structured score $\Delta(\mathbf{x},\mathbf{x}^{t})$, defined in Equation 8, aims to enforce a margin for each incorrectly identified edge.
(8) $\Delta(\mathbf{x},\mathbf{x}^{t})=\sum_{i=1}^{N}\sum_{j=1}^{N}\max(0,x_{ij}-x_{ij}^{t})\text{.}$
This approach eliminates the need to manually define the target adjacency matrix. Instead, it employs the output from the black-box optimiser to directly guide the updates of the model parameters. This methodology not only simplifies the process but also ensures a closer alignment with the solver’s procedural dynamics. To accelerate convergence, we integrate a hybrid of local and global loss. Relying solely on the global loss requires significantly more epochs to converge. Therefore, during training, we alternate batches between the global loss and the local loss, ensuring faster and more efficient model convergence.
5. Experimental Setup
We evaluate TSPRank’s effectiveness as a prediction layer integrated into various backbone models. Our goal is to demonstrate that TSPRank enhances performance compared to the original backbone and other general ranking algorithms across different domains. While achieving state-of-the-art (SOTA) performance still depends on the backbone design, TSPRank serves as a component to improve ranking performance. We focus on scenarios where (i) the data contain ordinal ranking labels rather than binary or relevance level labels, and (ii) the complete ranking or at least the top-k entities are important, extending the task’s interest beyond merely identifying the top-1 entity. We test our method with three datasets and backbone models from diverse domains: stock ranking, information retrieval, and historical events ordering. These datasets cover various data modalities, including tabular and textual data. All three tasks can be formulated as presented in Section 3.1. We aim to answer the following research questions (RQs):
  • RQ1: Does the pairwise-listwise TSPRank outperform the current SOTA general pairwise and listwise ranking methods across different backbone models and datasets?
  • RQ2: Does global learning lead to better performance?
  • RQ3: Does the pure pairwise method really consistently outperform the deep learning based listwise method across different domains?
  • RQ4: What are the advantages of hybrid pairwise-listwise TSPRank over other methods?
This section details the experimental setup for each task. All experiments were conducted on a single Nvidia A100 80G GPU.
5.1. Benchmark Models
We benchmark our pairwise-listwise TSPRank model against the original backbone models (if they originally contain a prediction layer), LambdaMART (Burges, 2010), and Rankformer (Buyl et al., 2023), as LambdaMART is still considered the SOTA pairwise ranking method and Rankformer is the SOTA deep learning-based listwise ranking method. The benchmark models chosen include the best pure pairwise and pure listwise methods, which is sufficient to demonstrate the effectiveness of our approach.
In practice, we replace the original prediction layer, which predicts the ranking scores for each entity, with different benchmark models to evaluate the performance achieved by this modification.
5.2. Datasets
Stock Ranking
The task of stock ranking focuses on accurately predicting and ranking stocks according to their anticipated future returns ratio, which aids investors in selecting stocks for investment purposes (Feng et al., 2019; Saha et al., 2021; Zhang et al., 2022). We use the dataset [^1] https://github.com/fulifeng/Temporal_Relational_Stock_Ranking/ introduced by Feng et al. (Feng et al., 2019), which includes historical trading data from 2013 to 2017 for two significant markets, NASDAQ and NYSE, containing 1026 and 1737 stocks, respectively. We maintain a consistent setting for splitting the data into training, validation, and testing sets over a 3-year, 1-year, and 1-year period.
Information Retrieval
Information retrieval is another critical area where ranking models are extensively applied. The task of information retrieval focuses on accurately ranking documents based on their relevance to a given query, which is crucial for search engines. LambdaMART and Rankformer are both originally proposed for information retrieval task (Burges, 2010; Buyl et al., 2023). We use the MQ2008-list [^2] https://www.microsoft.com/en-us/research/project/letor-learning-rank-information-retrieval/letor-4-0/ dataset from Microsoft LETOR4.0 (Qin and Liu, 2013), a popular benchmark dataset for LETOR algorithms. We acknowledge the existence of more recent information retrieval datasets, but they only provide relevance-level labels without ordinal ranking labels, making them unsuitable for our model evaluation.
Historical Events Ordering
In the final task, we include a textual dataset called “On This Day 2” (OTD2) [^3] https://github.com/ltorroba/machine-reading-historical-events for chronologically ordering historical events. Originally, Honovich et al. constructed this dataset for a regression task to predict the year of occurrence (Honovich et al., 2020). However, it can also be approached as a ranking task to predict the chronological order of events within a given group of events. The OTD2 dataset, sourced from the “On This Day” website, includes 71,484 events enriched with additional contextual information. Scraped in April 2020, it incorporates recent events and corrections by the site. Preprocessing excluded events before 1 CE and future projections such as “31st predicted perihelion passage of Halley’s Comet” in 2061 CE. We follow the 80%/10%/10% train/validation/test split as provided.
5.3. Backbones
Stock Ranking
We use the method by Feng et al. (Feng et al., 2019) as the backbone model, replacing the prediction layer with different benchmark ranking models. Although other potential methods such as STHAN-SR (Sawhney et al., 2021), ALSP-TF (Wang et al., 2022), and CI-STHPAN (Xia et al., 2024) have been proposed for stock selection, we do not include them due to the lack of last-state embeddings needed to integrate TSPRank, unavailability of source code, or reproducibility issues. Since our primary goal is to demonstrate TSPRank as a versatile ranking approach rather than achieving SOTA results on a single dataset, this selection of backbone is sufficient to validate our model’s performance.
Information Retrieval
Given that the features in the MQ2008-list are encoded using a combination of BM25 and TF-IDF methods, we treat BM25 and TF-IDF as the backbone models and use these encoded features as embeddings.
Historical Events Ordering
Concatenating event titles and information can result in contexts spanning thousands of tokens. Therefore, we use the text-embedding-3-small [^4] https://platform.openai.com/docs/guides/embeddings embeddings from OpenAI as the backbone model, which supports up to 8191 tokens. Our goal is not to achieve SOTA performance but to ensure a fair comparison. The small version (text-embedding-3-small) with 1536 dimensions, compared to the large version’s 3072 dimensions (text-embedding-3-large), is more computationally efficient and sufficient for our purposes.
5.4. Technical Setup
Stock Ranking
Considering that TSPRank specialises in small-scale ranking problems, we implement rankings within individual sectors, categorising stocks into groups based on their sectors. We exclude all sector groups containing three or fewer stocks and remove stocks with unidentified sector information. Detailed descriptions of the sector-specific grouping are provided in Appendix E. For the benchmarking process, we use outputs from the relational embedding layer introduced by Feng et al. (Feng et al., 2019), replacing the original prediction layer with the other benchmark models. This modification enables us to assess the extent to which transitioning from a point-wise (original model) to pairwise, listwise, and our hybrid approaches can enhance performance. All benchmark models and TSPRank with local learning are trained for 100 epochs for each sector, as 100 epochs are sufficient for the benchmark models to converge. We find that TSPRank with global learning generally requires more epochs to converge; therefore, TSPRank-global is trained for an additional 50 epochs to ensure convergence. Other hyperparameters are specified in Appendix B. For LambdaMART, we use 10,000 trees since previous experiments indicated that Rankformer failed to outperform LambdaMART with this configuration (Buyl et al., 2023). We choose the widely-used XGBoost [^5] https://xgboost.readthedocs.io/en/latest/tutorials/learning_to_rank.html (Chen and Guestrin, 2016) implementation of LambdaMART, especially due to the performance enhancements in its version 2.0.
Information Retrieval
In the MQ2008-list dataset, a query might contain over 1,000 documents, resulting in a graph that is too large for the TSP solver to handle efficiently. Therefore, we focus on a reranking or post-reranking stage. Specifically, we extract the top 10 and top 30 documents for each query to conduct two separate experiments. This approach allows us to manage the computational complexity while still evaluating the effectiveness of TSPRank in a realistic ranking scenario. We use similar hyperparameters as in the stock ranking task but increase the number of transformer layers in both Rankformer and TSPRank to four, as we do not use any pretrained embeddings here. To ensure fairness, we also follow the 5-fold cross-validation setup as provided by Microsoft.
Historical Events Ordering
We randomly allocate the events into ranking groups and train the ranking models to predict the ground truth order of occurrences within these groups. We test with group sizes of 10, 30 and 50 for fair comparisons. For each group size, we generate the group allocations using five different random seeds and report the average performance to mitigate the effects of random group allocation. Regarding the models, we observed overfitting with Rankformer and TSPRank when including the transformer block, even with a 0.5 dropout rate. This is likely due to the strong representational power of the contextual embedding and the relatively small size of the OTD2 dataset compared to other NLP corpora. Consequently, we retained only the scoring network in Rankformer and the trainable bilinear model in TSPRank, integrating them directly with the text-embedding-3-small embeddings. This setting not only prevents overfitting but also ensures similar model sizes. Other hyperparameters are specified in Appendix B. We remove the weights in the weighted loss function for TSPRank-Local here since all events should be weighted equally without any additional requirements in this task.
5.5. Evaluation
Stock Ranking
Given that our task encompasses both ranking and stock selection functionalities, we employ both traditional ranking metrics and financial performance metrics in our evaluation. For ranking effectiveness, we use the Mean Average Precision at $K$ (MAP@K), which evaluates the precision of the top $K$ ranked stocks by averaging the precision scores at each relevant stock position. We also use Kendall’s Tau, defined as: $\tau=\frac{(C-D)}{\frac{1}{2}n(n-1)}$, where $C$ is the number of concordant pairs (pairs of entities that are in the same order in both predicted and ground truth rankings), $D$ is the number of discordant pairs (pairs of entities that are in different orders in the rankings), and $n$ is the total number of entities being ranked. Kendall’s Tau is a correlation coefficient that measures the ordinal association between two ranked lists. This is suitable in our case, where we need to assess the similarity between the predicted rankings and the ground-truth rankings. Furthermore, to assess the financial impact of the rankings, we simulate trading activities by predicting the rankings of all individual stocks in every sector for the next trading day $t+1$ on trading day $t$. This simulation is consistent with the one presented by Feng et al. (Feng et al., 2019). We then rank these stocks, hold the top $k$ stocks $S_{k}$, and sell them on trading day $t+1$. The price of stock $i$ on day $t$ is denoted as $p_{i}^{(t)}$ per share. The cumulative Investment Return Ratio (IRR) and Sharpe Ratio (SR) are defined in Equation 9, where $R_{p}$ is the return of the portfolio, $R_{f}$ is the risk-free rate, and $\sigma$ is the standard deviation:
(9) $\text{IRR}^{k}=\sum_{i\in S^{k}}\frac{p_{i}^{(t)}-p_{i}^{(t-1)}}{p_{i}^{(t-1)}},\text{ SR}=\frac{E(R_{p}-R_{f})}{\sigma(R_{p})}.$
Information Retrieval
We evaluate the performance of our models using several well-established ranking metrics: Normalized Discounted Cumulative Gain (NDCG) at various cutoff levels (3, 5, 10), Mean Reciprocal Rank (MRR), and Kendall’s tau. NDCG measures a document’s usefulness based on its position in the result list, with higher positions receiving more weight. On the other hand, MRR assesses the rank position of the first relevant document, calculated as the average of the reciprocal ranks of the first relevant document for each query. The formula for MRR is defined by $\text{MRR}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\frac{1}{\text{rank}_{i}}$, where $|Q|$ is the total number of queries and $\text{rank}_{i}$ is the rank position of the first relevant document for the $i$-th query.
Historical Events Ordering
We report the Kendall’s Tau, MRR, Exact Match (EM) of ordering, and the Root Mean Square Error (RMSE) of the predicted and ground truth rankings. EM is defined as the ratio of the number of events whose predicted order exactly matches the ground truth order to the total number of events, formulated as: EM=# correctly ordered events# $\text{EM}=\frac{\text{\# correctly ordered events}}{\text{\# events}}$.
6. Results
Market Model $\tau$ IRR@1 SR@1 MAP@1 IRR@3 SR@3 MAP@3 IRR@5 SR@5 MAP@5
NASDAQ Feng et al. + MLP (Original) 0.0093 0.1947 0.5341 0.1690 0.2366 0.9881 0.3253 0.1892 0.9682 0.5871
Feng et al. + LambdaMART 0.0071 0.0310 -0.0873 0.1539 0.0340 0.0445 0.3144 0.0505 0.2678 0.5858
Feng et al. + Rankformer 0.0110 0.2257 0.5464 0.1620 0.2857 1.1245 0.3216 0.2309 1.0943 0.5860
Feng et al. + TSPRank-Local 0.0291 0.5353 1.2858 0.1658 0.4416 1.7401 0.3297 0.2537 1.2623 0.5932
Feng et al. + TSPRank-Global 0.0447 0.7849 1.7471 0.1633 0.5224 2.0359 0.3364 0.2937 1.4331 0.5999
NYSE Feng et al. + MLP (Original) 0.0162 0.4170 1.0755 0.1791 0.2574 1.2367 0.2841 0.2257 1.3186 0.4649
Feng et al. + LambdaMART 0.0054 0.1005 0.1367 0.1307 0.0732 0.4192 0.2592 0.1063 0.6882 0.4574
Feng et al. + Rankformer 0.0181 0.2924 0.9113 0.1535 0.2701 1.2890 0.2758 0.2515 1.4200 0.4651
Feng et al. + TSPRank-Local 0.0313 0.5012 1.5710 0.1424 0.3974 1.9735 0.2756 0.2788 1.6662 0.4680
Feng et al. + TSPRank-Global 0.0422 0.4787 1.4552 0.1392 0.3889 1.9976 0.2756 0.2816 1.7350 0.4732
Table 1. Performance comparison of Feng et al., LambdaMART, Rankformer, and TSPRank on the NASDAQ and NYSE stock ranking dataset, averaged across all filtered sectors.
Top 10 Top 30
Model Type NDCG@3 NDCG@5 NDCG@10 MRR $\tau$ NDCG@3 NDCG@5 NDCG@10 MRR $\tau$
LambdaMART Pairwise 0.6833 0.7222 0.8707 0.4259 0.1474 0.7340 0.7298 0.7403 0.3617 0.2372
Rankformer Listwise 0.7220 0.7565 0.8865 0.4661 0.2317 0.7486 0.7470 0.7596 0.3732 0.2834
TSPRank-Local Pairwise-Listwise 0.6858 0.7213 0.8719 0.4266 0.1544 0.7189 0.7240 0.7362 0.3206 0.2054
TSPRank-Global Pairwise-Listwise 0.7281 0.7585 0.8884 0.4861 0.2212 0.7582 0.7558 0.7631 0.3895 0.2647
Table 2. Evaluation of LambdaMART, Rankformer, and TSPRank on MQ2008-list information retrieval dataset for top 10 and top 30 documents.
Group Size 10 30 50
Model Type $\tau$uparrow EM $\uparrow$ MRR $\uparrow$ RMSE $\downarrow$ $\tau$uparrow EM $\uparrow$ MRR $\uparrow$ RMSE $\downarrow$ $\tau$uparrow EM $\uparrow$ MRR $\uparrow$ RMSE $\downarrow$
te-3-small + LambdaMART Pairwise 0.6297 0.3008 0.7554 1.993 0.5929 0.1064 0.6122 5.969 0.6000 0.0639 0.5596 9.618
te-3-small + Rankformer Listwise 0.6190 0.2899 0.7361 1.998 0.5859 0.0921 0.4911 5.973 0.5724 0.0527 0.3526 10.069
te-3-small + TSPRank-Local Pairwise-Listwise 0.5658 0.2856 0.7679 2.296 0.5095 0.0873 0.5739 6.930 0.4713 0.0460 0.3949 12.084
te-3-small + TSPRank-Global Pairwise-Listwise 0.6301 0.3350 0.7936 2.057 0.6302 0.1384 0.7300 5.770 0.6207 0.0871 0.6618 9.602
Table 3. Evaluation of LambdaMART, Rankformer, and TSPRank on OTD2 dataset for historical events ordering for group sizes of 10, 30, and 50. “te-3-small” stands for “text-embedding-3-small”.
The aggregated results for all benchmark models on the stock ranking dataset are presented in Table 1. These results represent the average performance across 50 sectors in the NASDAQ market and 70 sectors in the NYSE market, excluding sectors with three or fewer stocks as outlined in the experimental setup. Detailed results for each individual sector are provided in Appendix E. The results on the MQ2008-list retrieval dataset are shown in Table 2. A 5-fold cross-validation was performed, with detailed results for each fold provided in Appendix F. Lastly, the results on OTD2 for historical events ordering are presented in Table 3, with complete results for different random group allocations provided in Appendix G.
The results highlight the comparisons among three types of models: pairwise (LambdaMART), listwise (Rankformer), and our pairwise-listwise method (TSPRank).
Better Performance of Pairwise-Listwise Method Across Diverse Tasks (RQ1). Our pairwise-listwise method, TSPRank-Global, demonstrates outstanding robustness and superior performance across diverse datasets and domains. For instance, on the NASDAQ stock ranking dataset (Table 1), TSPRank-Global achieves a Kendall’s Tau of 0.0447, significantly surpassing both LambdaMART (0.0071) and Rankformer (0.0110). Similarly, on the NYSE dataset (Table 1), TSPRank-Global attains a Kendall’s Tau of 0.0422, outperforming LambdaMART (0.0054) and Rankformer (0.0181). Both TSPRank-Local and TSPRank-Global also consistently achieve higher IRR and SR, yielding improved financial performance. The model’s robustness is further evidenced by its consistent top performance across other datasets. For example, in the MQ2008-list (Table 2), TSPRank-Global leads with NDCG@10 scores of 0.8884 for the top 10 documents and 0.7631 for the top 30 documents, indicating superior ranking accuracy, despite a minor difference in Kendall’s Tau compared to Rankformer (around 0.01 to 0.02). Additionally, TSPRank-Global achieves top performance in the historical events ordering task (Table 3), regardless of the group size (10, 30, and 50).
TSPRank Can Be Extra Component Upon Embeddings For Boosting The Performance. In the stock ranking and ordering tasks for historical events, we directly deployed TSPRank on the stock embeddings and text embeddings. As demonstrated, TSPRank effectively serves as an additional component on embeddings, significantly enhancing ranking performance. This capability allows TSPRank to have broader applications across various domains.
Stronger Robustness Towards Number of entities. TSPRank-Global also shows greater robustness concerning the number of entities compared to the pure pairwise LambdaMART and listwise Rankformer. As evidenced by the MQ2008 (Table 2) results, the gap between LambdaMART and Rankformer narrows as the number of documents increases from 10 to 30, with the NDCG@3 score difference between these two methods reducing from -0.0387 to -0.0146. Similarly, in historical events ordering (Table 3), LambdaMART and Rankformer’s performance fluctuates more significantly with group size changes. This is evident from the varying differences in their Kendall’s Tau and MRR compared to those of TSPRank-Global. Conversely, TSPRank-Global consistently performs well regardless of group size, maintaining top performance across most metrics.
Global Learning Outperforms Local Learning (RQ2). As evidenced by the performance of TSPRank-Global compared to TSPRank-Local on MQ2008 and OTD2. For example, on the OTD2 dataset (Table 3), TSPRank-Global achieves a Kendall’s Tau of 0.6301, 0.6302, 0.6207 for the three different group sizes, compared to TSPRank-Local’s 0.5658, 0.5095, and 0.4713. These results suggest that the end-to-end global optimisation approach of TSPRank-Global aligns more effectively with overall optimisation goals, whereas the non-end-to-end method may not perfectly align with global optimisation by the TSP solver.
LambdaMART Does Not Always Outperform Deep Learning Based Ranking Algorithm (RQ3). Although Qin et al. suggests that GBDT models, such as LambdaMART, generally outperform deep learning models on standard information retrieval datasets (Qin et al., 2021), our results indicate that this is not always the case in other contexts. As demonstrated in Table 1, within the domain-specific scenario of stock ranking, LambdaMART shows markedly poorer performance. Specifically, LambdaMART achieves a Kendall’s Tau of 0.0071 on NASDAQ and 0.0054 on NYSE, trailing behind Rankformer (0.0110 and 0.0181) and TSPRank-Global (0.0447 and 0.0422). Table 2 further validates this observation. In both the top 10 and top 30 document settings, LambdaMART consistently underperforms Rankformer and TSPRank-Global when adapting to ordinal ranking labels across all metrics. These results underscore that while LambdaMART performs well with binary relevance labels or relevance levels labels, it is not always the optimal choice for recovering complete rankings.
Overall, TSPRank’s success suggests that combining pairwise and listwise approaches helps the model capture the relative ordering of entities and the overall ranking structure. This dual focus likely contributes to its superior performance across various datasets and metrics, making it versatile and powerful.
7. Visualisation Analysis
As mentioned before, our pairwise-listwise TSPRank addresses both the lack of listwise optimisation in pairwise LambdaMART and the robustness issues in listwise Rankformer, which arise from attempting to predict the complete rankings directly.
To empirically validate this and answer RQ4 (the advantages of hybrid pairwise-listwise TSPRank over other methods), we conducted another experiment on the OTD2 dataset, chosen because of the better interpretability provided by text. We arbitrarily sample three events each from the US, UK, and China, ensuring that the events within each country occurred at least 50 years apart (detailed in Table 4). We then use the pretrained LambdaMART, Rankformer, and TSPRank-Global from the experiments described in Section 5 to predict the rankings for this constructed group. We visualise the predictions given by the three rankers in Figure 2. Both the pairwise LambdaMART and pairwise-listwise TSPRank-Global successfully recover the intra-country event order, correctly identifying the relative order within the same colour group. However, the listwise Rankformer incorrectly predicts that “US-2” occurs earlier than “US-1”.
Although LambdaMART successfully recovers the intra-country event order, it makes mistakes in the overall ranking, particularly failing to recover the correct order of [“US-1”, “UK-1”, “CN-1”] and [“US-3”, “UK-3”, “CN-3”], validating its weaker ability for listwise ranking optimisation. In contrast, TSPRank accurately recovers most orders, with only a minor error between “US-2” and “UK-2”, which only have a 30-year gap (1934 and 1904).
[…]
Event Title Year Rank Label
1st US store to install electric lights, Philadelphia 1878 3 US-1
1st sitting US President to visit South America, FDR in Colombia 1934 5 US-2
75th US Masters Tournament, Augusta National GC: Charl Schwartzel of South Africa birdies the final 4 holes to win his first major title, 2 strokes ahead of Australian pair Adam Scott and Jason Day 2011 8 US-3
Charles Watson-Wentworth, 2nd Marquess of Rockingham, becomes Prime Minister of Great Britain 1782 2 UK-1
1st main line electric train in UK (Liverpool to Southport) 1904 4 UK-2
UK Terrorism Act 2006 becomes law 2006 7 UK-3
A Mongolian victory at the naval Battle of Yamen ends the Song Dynasty in China 1279 1 CN-1
US Senate rejects China People’s Republic membership to UN 1953 6 CN-2
China’s Hubei province, the original center of the coronavirus COVID-19 outbreak eases restrictions on travel after a nearly two-month lockdown 2020 9 CN-3
Table 4. Event titles in the constructed group. Labels indicate the order of occurrence within each country, e.g., “US-1” denotes the earliest event in the US within the group.
Figure 2. Visualisation of predictions by LambdaMART, Rankformer, and TSPRank-Global on the constructed group. Numbers in parentheses indicate the true ranking.
Figure 3. Illustration of the intra-country pairwise comparison graph. Edges between pairs of events from different countries are omitted for clarity. Scores highlighted in red indicate errors in the pairwise prediction for TSPRank-Global.
8. Inference Latency
Incorporating a combinatorial optimisation solver is known to cause additional computational overhead, which can impact the inference latency. Therefore, we compare the average inference time per ranking group for TSPRank and Rankformer on the OTD2 historical events ordering dataset. The experimental setup and detailed results are provided in Appendix D.
We observe that the inference time for TSPRank increases exponentially with the number of ranking entities, primarily due to the discrete TSP solver. When the group size reaches 100, the solver consumes approximately 99.8% of the inference time, while the encoding time remains similar to that of Rankformer. Therefore, TSPRank is more suitable for small-scale ranking problems with fewer than 30 entities, where its inference time is comparable to Rankformer. It is also advantageous in scenarios where TSPRank’s performance improvements outweigh the increased latency. Solvers like Gurobi are efficient for small TSP graphs. However, exact solutions for larger graphs require specialised techniques, such as the Concorde TSP Solver [^6] https://www.math.uwaterloo.ca/tsp/concorde/index.html, which uses cutting-plane algorithms and a branch-and-bound approach to solve graphs with up to 85,900 nodes (Applegate, 2006). Beyond exact solvers, heuristic algorithms like the Christofides–Serdyukov algorithm (Christofides, 2022) and the Lin–Kernighan–Helsgaun heuristic (Lin and Kernighan, 1973; Helsgaun, 2000, 2017) are also used.
Moreover, recent advances in neural networks have enabled the use of deep learning to approximate the TSP reward function, employing reinforcement learning to solve the problem (Bello et al., 2017; Khalil et al., 2017; Kool et al., 2019). While we do not delve deeply into the various TSP solutions in this paper, we demonstrate that inference latency stems from the TSP solver and discuss potential enhancements for reducing inference latency and tackling large-scale ranking problems.
9. Conclusion
This study introduces TSPRank, a novel pairwise-listwise approach for position-based ranking tasks, by modelling them as the Travelling Salesman Problem (TSP). We present two learning methods for TSPRank, integrating listwise optimization into pairwise comparisons to address the limitations of traditional pure pairwise and listwise ranking models. TSPRank’s application to diverse backbone models and datasets, including stock ranking, information retrieval, and historical events ordering, demonstrates its superior performance and robustness. Key findings include TSPRank’s ability to outperform existing models such as LambdaMART and Rankformer across multiple datasets and various metrics. Furthermore, TSPRank’s capability to function as an additional component on embeddings suggests its versatility and potential for broader applications across different domains.
Our empirical analysis reveals that TSPRank’s superior performance is due to the discrete TSP solver’s ability to more effectively utilise listwise information. It also enhances TSPRank’s tolerance to uncertainties in pairwise comparisons by ensuring the selection of the highest-scoring permutation. Future work could explore the use of alternative TSP solvers and further enhancements to the Gurobi solver to reduce inference latency and improve scalability.
Appendix A TSP Optimisation Problem
Ranking entity $\mathbf{e}_{j}$ immediately after entity $\mathbf{e}_{i}$ is analogous to travelling from city $i$ to city $j$. Let $N$ be the total number of entities to be ranked and $z_{i}$ variables to represent the number of entities ranked before entity $i$. The MILP formulation of the TSPRank inference is as follows:
(10) $\min\limits_{x_{ij}}$ $\quad\sum\limits_{i=1}^{N}\sum\limits_{j=1,j\neq i}^{N}a_{ij}x_{ij}$
(11) $s.t.$ $\quad\sum_{j=1,j\neq i}^{N}x_{ij}\leq 1\quad\text{ for all }i$
(12) $\quad\sum_{i=1,i\neq j}^{N}x_{ij}\leq 1\quad\text{ for all }j$
(13) $\quad\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}x_{ij}=N-1$
(14) $\quad z_{i}+1\leq z_{j}+N(1-x_{ij})\quad i,j=2,\ldots,N,\;i\neq j$
(15) $\quad z_{i}\geq 0\quad i=2,\ldots,N$
Constraints (11) and (12) ensure that each entity has at most one predecessor and one successor in the ranking, expressed as $22N-$ 2 linear inequalities. Constraint (13) ensures that the total number of pairwise comparisons is exactly $N-1$. These constraints ensure that the selected set of pairwise comparisons forms a valid ranking sequence. Constraints (14) and (15) introduce variables $z$ to eliminate multiple separate sequences and enforce that there is a single, complete ranking that includes all entities.
Appendix B Hyperparameters Setting
The hyperparameters are specified in Table 5. In the table, several abbreviations are used for conciseness and clarity:
  • lr: Learning rate, which controls the step size during the optimization process.
  • tf layer: Number of transformer layers, indicating the depth of the transformer model.
  • tf nheads: Number of attention heads, which specifies the number of parallel attention mechanisms within the transformer layer.
  • tf dim_ff: Dimensionality of the feed-forward network within the transformer layer.
Dataset Hyperparameter Value Model
All Tasks n_estimators $10^{4}$ LambdaMART
loss_fn rank:pairwise
eval_metric auc
early_stopping 50
loss OrdinalLoss (Buyl et al., 2023) Rankformer
Stocks lr 1e-4 Rankformer, TSPRank
weight_decay 1e-5
tf layer 1
tf nheads 8
tf dim_ff 128
batch_size 128
MQ2008 lr 1e-4 Rankformer, TSPRank
weight_decay 1e-5
tf layer 4
tf nheads 8
tf dim_ff 128
batch size 64
Historical Events lr 1e-4 Rankformer, TSPRank
weight_decay 1e-5
tf dim_ff 0
batch size 32
Table 5. Hyperparameters for LambdaMART model across different tasks: Stock Ranking, MQ2008, and Historical Events Ordering.
Appendix C Additional Visualisation Examples
Figure 4. Visualisation of the model predictions on additional groups. Numbers in parentheses indicate the true ranking.
We further construct three more groups using events from different countries. The event titles, labels, and true rankings are provided in Table 6. The visualisation of the models’ predictions are shown in Figure 4. We observe similar patterns to those discussed in Section 7.
Group Event Title Year Rank Label
1 Japanese immigration to Brazil begins when 781 people arrive in Santos aboard the Kasato-Maru ship 1908 4 JP-1
Passenger ship Lurline sends radio signal of sighting Japanese war fleet 1941 5 JP-2
A reactor at the Fukushima Daiichi nuclear power plant melts and explodes and releases radioactivity into the atmosphere a day after Japan’s earthquake. 2011 8 JP-3
The State Bank of the Russian Empire is established. 1860 3 RU-1
”Mirror”, Russian film directed by Andrei Tarkovsky, starring Margarita Terekhova and Ignat Daniltsev, is released 1975 7 RU-2
Russian city Moscow begins a city-wide lockdown after 4 hours notice due to coronavirus COVID-19 2020 9 RU-3
Albrecht II of Habsburg becomes king of Germany 1438 1 DE-1
Magdeburg in Germany seized by forces of the Holy Roman Empire under earl Johann Tilly, most inhabitants massacred, one of the bloodiest incidents of the Thirty Years’ War 1631 2 DE-2
The cult classic ”One Million Years B.C.”, starring Raquel Welch, is released 1st in West Germany 1966 6 DE-3
2 French King Charles VIII occupies Florence 1494 2 FR-1
1st steamboat, Pyroscaphe, 1st run in France 1783 3 FR-2
Francois Mitterrand becomes president of France 1981 7 FR-3
The Hangul alphabet is published in Korea 1446 1 KR-1
Korea is divided into North and South Korea along the 38th parallel 1945 6 KR-2
North Korea blocks a South Korean supply delegation from the Kaesong joint industrial zone 2013 8 KR-3
Bahia Independence Day: the end of Portuguese rule in Brazil, with the final defeat of the Portuguese crown loyalists in the province of Bahia 1823 4 BR-1
Belo Horizonte, the first planned city of Brazil, founded 1897 5 BR-2
Brazilian court blocks President Michel Temer from abolishing Renca, which would open parts of the Amazon to mining 2017 9 BR-3
3 Indian Mutiny against rule by the British East India Company begins with the revolt of the Sepoy soldiers in Meerut 1857 1 IN-1
Gandhi supports the African People’s Organisations resolution to declare the Prince of Wales day of arrival in South Africa a day of mourning, in protest against the South Africa Acts disenfranchisement of Indians, Coloureds and Africans 1910 4 IN-2
Bomb attack on train in Assam India (27 soldiers killed) 1995 7 IN-3
1st Italian Parliament meets at Turin 1860 2 IT-1
King Victor Emmanuel III of Italy abdicates and is succeeded by his son Umberto II who reigns for only 34 days before the monarchy is abolished 1946 5 IT-2
Storms in Italy kill at least 11 with 75% of Venice flooded and two tornadoes striking Terracina 2018 9 IT-3
Edmund Barton is elected Prime Minister in Australia’s first parliamentary election 1901 3 AU-1
Australian Championships Women’s Tennis: Beryl Penrose wins her only Australian singles title; beats Thelma Coyne Long 6-4, 6-3 1955 6 AU-2
Cricket World Cup, Melbourne (MCG): Australia defeats fellow host New Zealand by 7 wickets to win their 5th title; Player of Series: Mitchel Starc 2015 8 AU-3
Table 6. Three additionally constructed groups using events from different countries.
Appendix D Latency Analysis Details
We conduct an inference time analysis on a standalone device with an AMD Ryzen 7 5800X 8-Core CPU and an Nvidia RTX 4070 Super GPU. Using the standalone device instead of the one used for training is to avoid interference from other running jobs on the shared computing node.
Figure 5 illustrates the average inference time in seconds per ranking group for TSPRank and Rankformer on the OTD2 historical events ordering dataset, with an input dimensionality of 1536. We provide the inference times for various group sizes: 5, 10, 30, 50, and 100.
The comparison shows that the inference time for TSPRank increases exponentially with the number of ranking entities in a group. This increase is primarily due to the discrete TSP solver, as indicated by the overlapping blue and red lines. The blue line represents the time taken by the TSP solver, while the red line shows the total inference time for a ranking group. When the group size reaches 100, approximately 99.8% of the inference time is consumed by the solver. The encoding time from the input to the adjacency matrix remains similar to that of Rankformer (purple).
Figure 5. Mean inference time per ranking group for TSPRank and Rankformer on OTD2 across group sizes of 5, 10, 30, 50, and 100. The blue (TSPRank solver only) and red (TSPRank) lines overlap with minor differences.
Appendix E Complete Stocks Ranking Results For Sectors
Table 7 to 16 presents the comprehensive results for each sector in NASDAQ and NYSE after filtering.
Sector Stocks $\tau$ IRR@1 SR@1 MAP@1 IRR@3 SR@3 MAP@3 IRR@5 SR@5 MAP@5
Computer Manufacturing 5 -0.0521 -0.3767 -1.2666 0.1857 0.1037 0.5894 0.4529 0.2167 1.3724 1.0000
Transportation Services 5 -0.0185 -0.3402 -1.1114 0.1941 0.2663 1.2239 0.4590 0.1263 0.7757 1.0000
Medical/Dental Instruments 17 0.0227 -0.1845 -0.2096 0.1055 0.2604 1.2060 0.1519 0.3940 2.0538 0.1922
Property-Casualty Insurers 19 0.0044 -0.3242 -0.5479 0.1055 -0.1942 -0.4679 0.2075 -0.2643 -1.0239 0.2131
Medical Specialities 5 -0.0101 -0.0157 0.0948 0.1814 0.0702 0.4433 0.4902 -0.0402 -0.1500 1.0000
Biotechnology: Biological Products (No Diagnostic Substances) 12 0.0288 0.5664 1.1922 0.1224 0.7839 2.1772 0.2253 0.4882 1.7150 0.3138
Telecommunications Equipment 16 -0.0046 -0.6169 -1.1101 0.1646 -0.2057 -0.4332 0.1718 -0.1449 -0.4478 0.2132
Medical/Nursing Services 6 0.0507 1.0367 1.8435 0.1688 0.7042 2.3042 0.4062 0.4471 1.8070 0.7530
Water Supply 9 0.0131 0.2849 1.0457 0.1561 0.2404 1.1933 0.2417 0.2146 1.1336 0.4176
Real Estate Investment Trusts 8 -0.0084 -0.0165 -0.0359 0.1055 0.0948 0.8628 0.2625 0.0099 0.1480 0.4738
Business Services 17 0.0130 0.4041 1.1789 0.1266 0.2231 1.1491 0.1589 0.3422 1.7452 0.1987
Major Pharmaceuticals 29 0.0008 0.2889 0.7399 0.0970 0.7151 1.4622 0.1273 0.4020 1.2343 0.1419
Farming/Seeds/Milling 6 0.0227 0.1512 0.7310 0.1688 -0.0104 0.0358 0.3945 0.0634 0.4711 0.7495
Industrial Specialties 11 0.0346 2.0118 4.1591 0.1224 0.7883 2.8012 0.2304 0.4046 2.0223 0.3336
Air Freight/Delivery Services 6 -0.0014 -0.0397 0.0735 0.1899 -0.0147 0.0708 0.3680 0.0704 0.4319 0.7407
Electrical Products 13 -0.0202 -0.2313 -0.6316 0.0675 -0.0352 -0.0250 0.1517 0.0833 0.5303 0.2264
Television Services 8 -0.0054 -0.0664 -0.0262 0.1224 -0.1397 -0.4674 0.2876 -0.0528 -0.1493 0.4848
Investment Bankers/Brokers/Service 18 0.0195 1.5618 3.0264 0.1899 0.2890 1.8445 0.2187 0.2075 1.3520 0.2438
Catalog/Specialty Distribution 7 0.0210 -0.0168 0.1895 0.1814 0.5484 1.8405 0.3354 0.4288 1.8188 0.5961
Savings Institutions 20 0.0460 0.5287 1.5253 0.1181 0.3299 1.5116 0.1646 0.2436 1.2962 0.1981
Other Consumer Services 9 0.0079 -0.0556 0.3305 0.2532 -0.2156 -0.6957 0.2904 -0.0076 0.0782 0.4044
Auto Parts:O.E.M. 6 -0.0104 -0.1260 -0.1430 0.2152 -0.1427 -0.5708 0.3638 0.0074 0.1431 0.7229
Trucking Freight/Courier Services 11 0.0265 0.6905 1.5795 0.1772 0.4522 1.4994 0.2501 0.5216 1.8957 0.3417
Radio And TV Broadcasting And Communications Equipment 10 -0.0055 0.1535 0.5546 0.0970 0.0259 0.2365 0.2168 0.2805 1.2102 0.3377
Biotechnology: Electromedical & Electrotherapeutic Apparatus 5 0.0513 1.1212 2.2247 0.2532 0.8664 2.9716 0.5070 0.3760 1.9405 1.0000
Home Furnishings 8 0.0303 -0.1578 -0.3445 0.1308 0.1396 0.6085 0.2752 0.1764 0.8462 0.4762
Major Chemicals 7 -0.0118 -0.3026 -0.8089 0.1730 0.0614 0.3825 0.3291 0.0943 0.6049 0.5455
Other Specialty Stores 8 0.0027 0.4400 1.1201 0.1561 -0.1071 -0.3697 0.2623 0.0242 0.2259 0.4928
Computer Software: Programming, Data Processing 7 -0.0118 -0.1717 -0.2653 0.2363 0.4192 1.6818 0.3425 0.2852 1.5359 0.5278
Restaurants 20 -0.0264 -0.4583 -1.2790 0.0464 -0.0233 0.0328 0.1416 -0.0919 -0.3906 0.1653
Automotive Aftermarket 6 0.0059 0.0265 0.2472 0.1646 0.1734 0.9593 0.3872 0.2190 1.3142 0.7483
Investment Managers 5 0.0092 0.2662 1.3199 0.2236 0.1422 0.9601 0.4869 0.0976 0.7306 1.0000
Oil Refining/Marketing 6 -0.0132 -0.2807 -1.0433 0.2785 0.1007 0.5836 0.3687 0.1559 0.9894 0.6879
Real Estate 7 0.0226 0.3281 1.2869 0.1857 0.2733 1.6266 0.3202 0.2303 1.5590 0.6019
Apparel 8 -0.0081 0.0019 0.2242 0.1308 0.3005 1.1491 0.2600 0.1953 0.9464 0.4672
Oil & Gas Production 6 -0.0036 -0.1716 -0.3582 0.1730 -0.3052 -1.1493 0.3697 -0.2423 -0.8614 0.7456
Computer Communications Equipment 5 -0.0025 0.8757 1.9336 0.2152 0.2596 1.3220 0.4681 0.0383 0.3025 1.0000
Clothing/Shoe/Accessory Stores 8 -0.0153 -0.1938 -0.2218 0.1730 -0.2163 -0.4854 0.2675 -0.0573 -0.0205 0.4725
Commercial Banks 5 0.0034 0.0180 0.1952 0.1561 0.2602 1.4412 0.5117 0.1656 0.9570 1.0000
Hospital/Nursing Management 5 -0.0244 0.0170 0.3074 0.2278 0.1864 0.9844 0.4473 0.2216 1.3716 1.0000
Banks 5 0.0689 0.4518 1.8758 0.2194 0.3074 1.6594 0.5326 0.0939 0.6403 1.0000
Professional Services 6 0.0350 -0.2014 -0.6058 0.1646 0.2080 1.1435 0.3901 0.1859 1.1313 0.7501
Packaged Foods 7 0.0162 -0.0081 0.1270 0.1688 -0.0930 -0.4279 0.$3$0.0692 -0.3715 0.5875
Metal Fabrications 12 -0.0010 -0.1004 -0.1038 0.1392 0.1738 0.8350 0.2014 0.2037 1.0968 0.2687
Diversified Commercial Services 6 0.0294 0.0471 0.3382 0.2658 0.5017 2.1019 0.4004 0.2662 1.5863 0.7258
Computer peripheral equipment 6 0.0434 0.3923 1.0476 0.2110 0.2852 1.2401 0.4163 0.2432 1.3341 0.7431
Steel/Iron Ore 7 0.0250 0.6945 1.8428 0.1814 0.4356 1.7598 0.3472 0.1355 0.7523 0.5850
Hotels/Resorts 5 0.0471 0.9652 2.2147 0.2405 1.2521 4.3657 0.5136 0.8753 3.8558 1.0000
Biotechnology: In Vitro & In Vivo Diagnostic Substances 5 0.0126 0.3358 0.9351 0.1857 0.5339 1.9262 0.4754 0.4424 2.1666 1.0000
Electronic Components 8 0.0081 0.5326 1.3140 0.1350 0.7561 2.3124 0.2850 0.7520 2.7036 0.4646
Average 9.12 0.0094 0.1947 0.5341 0.1690 0.2366 0.9881 0.3254 0.1892 0.9682 0.5871
Table 7. Performance of baseline model proposed in (Feng et al., 2019) on NASDAQ
Sector Stocks $\tau$ IRR@1 SR@1 MAP@1 IRR@3 SR@3 MAP@3 IRR@5 SR@5 MAP@5
Biotechnology: Laboratory Analytical Instruments 5 0.0244 0.9165 2.6132 0.2447 0.5556 3.2469 0.4930 0.4553 3.1498 1.0000
Diversified Commercial Services 9 0.0315 1.9463 2.2132 0.2827 1.0483 2.8712 0.3033 0.7726 2.9765 0.3992
Other Specialty Stores 11 -0.0083 -0.2757 -0.4046 0.1350 0.1972 0.7898 0.2220 0.2008 0.9486 0.3136
Investment Managers 18 0.0147 1.2026 3.1766 0.1224 0.7153 3.3275 0.1547 0.5711 3.2101 0.2090
Electrical Products 13 0.0021 0.2232 0.7774 0.1266 0.1747 0.9553 0.1699 0.2038 1.2445 0.2505
Major Pharmaceuticals 18 0.0083 -0.1707 -0.0414 0.1730 -0.1365 -0.3281 0.2218 -0.0340 -0.0488 0.2218
Other Pharmaceuticals 5 0.0420 0.7478 2.6057 0.2700 0.2195 1.2131 0.4916 0.1719 1.0601 1.0000
Automotive Aftermarket 10 0.0289 0.2186 0.8903 0.0717 0.1810 0.7936 0.2117 0.1168 0.5851 0.3680
Precious Metals 11 -0.0163 0.6813 1.6358 0.1857 0.0364 0.2783 0.2150 0.0238 0.2233 0.3088
Publishing 5 0.0151 -0.4553 -1.1204 0.2743 -0.1287 -0.3477 0.4744 -0.1218 -0.5510 1.0000
Military/Government/Technical 11 0.0094 0.3709 1.3192 0.1013 0.2765 1.3624 0.1962 0.1733 1.0585 0.3280
Business Services 22 0.0371 0.8236 2.3041 0.0633 0.5385 2.6426 0.1423 0.6044 3.6581 0.1825
Packaged Foods 16 0.0130 0.0371 0.3576 0.1055 0.2376 1.1822 0.1892 0.2100 1.2396 0.2425
Life Insurance 17 0.0316 0.4408 1.1822 0.1730 0.3049 1.3894 0.1927 0.4008 2.1153 0.2161
Power Generation 15 -0.0011 -0.2409 -1.1783 0.1181 0.0223 0.2317 0.1723 0.0708 0.6655 0.2197
Clothing/Shoe/Accessory Stores 15 0.0175 1.9787 2.2021 0.1097 0.3421 0.9112 0.1711 0.1680 0.6053 0.2362
Medical Specialities 11 0.0526 0.2179 0.7375 0.2194 0.5787 2.5074 0.2525 0.4993 2.7790 0.3406
Property-Casualty Insurers 30 0.0192 0.2673 0.7250 0.1561 0.0009 0.1354 0.1425 0.0476 0.3467 0.1597
Accident &Health Insurance 5 -0.0143 0.4029 1.0810 0.2785 0.4239 2.0508 0.4489 0.2645 1.7726 1.0000
Farming/Seeds/Milling 7 0.0286 0.5901 2.1545 0.2110 0.0935 0.7004 0.3469 0.0952 0.7942 0.5747
Engineering & Construction 5 0.0126 -0.4365 -0.4714 0.3080 -0.0482 0.0147 0.4684 -0.0339 -0.0130 1.0000
Aerospace 9 0.0033 0.0075 0.2826 0.2110 0.2362 1.0729 0.2651 0.3670 1.9615 0.3881
Industrial Specialties 9 0.0000 1.0994 1.8806 0.2025 0.4923 1.9735 0.2921 0.3469 1.9804 0.3838
Specialty Insurers 6 0.0311 0.4101 2.1419 0.2236 0.3361 2.9795 0.3976 0.3112 3.1113 0.7424
Major Chemicals 29 0.0367 2.0708 1.9914 0.0591 0.8546 2.3885 0.1067 0.7252 2.8721 0.1438
Integrated oil Companies 19 0.0051 -0.1042 -0.0241 0.1181 0.2789 1.1388 0.1840 0.2697 1.2332 0.2238
Air Freight/Delivery Services 8 0.0255 -0.1557 0.0996 0.2152 0.2922 1.1280 0.3223 0.1947 1.0094 0.4667
Auto Parts:O.E.M. 18 0.0028 0.0070 0.1948 0.1308 0.0986 0.5461 0.1505 0.1639 0.8894 0.1828
Metal Fabrications 23 -0.0221 -0.2684 -0.7043 0.0802 -0.1461 -0.4471 0.1493 -0.0714 -0.2014 0.1765
Professional Services 11 0.0264 0.3941 1.4152 0.1181 0.4917 2.3416 0.1934 0.4289 2.3950 0.3339
Consumer Electronics/Appliances 6 0.0177 0.0551 0.3653 0.2152 0.2219 1.5692 0.3870 0.1651 1.4599 0.7282
Fluid Controls 7 0.0054 0.4329 1.2246 0.1899 0.0558 0.3530 0.3397 0.1384 0.7871 0.5731
Other Consumer Services 17 0.0035 0.2075 0.6701 0.1350 -0.0566 -0.0291 0.1894 0.0480 0.3380 0.2213
Steel/Iron Ore 9 0.0161 1.3380 1.9147 0.1730 0.7475 1.9947 0.2719 0.2903 1.1482 0.4063
Oil/Gas Transmission 6 0.0597 0.1326 0.7732 0.1941 0.2167 1.7503 0.4048 0.1632 1.5104 0.7235
Plastic Products 6 0.0434 0.5573 1.8102 0.2068 0.2234 1.1718 0.4102 0.2424 1.6670 0.7250
Agricultural Chemicals 7 -0.0338 -0.0122 0.0616 0.1308 -0.0179 0.0328 0.3043 0.0169 0.1883 0.5500
Containers/Packaging 9 0.0532 0.3567 1.6435 0.1857 0.4655 2.9107 0.2787 0.3857 2.8297 0.4273
Water Supply 7 0.0738 0.7212 1.9706 0.2489 0.6259 2.6691 0.3610 0.5509 2.8809 0.6004
Finance: Consumer Services 20 -0.0260 0.2587 0.9367 0.1097 0.3842 1.5753 0.1486 -0.0273 -0.0406 0.1818
Commercial Banks 21 0.0054 0.2928 1.1736 0.1055 0.4938 2.5922 0.1517 0.4329 2.4692 0.1838
Medical/Dental Instruments 12 0.0329 0.7236 2.3267 0.1477 0.7127 3.6677 0.2138 0.4851 3.0057 0.3029
Computer Software: Prepackaged Software 11 0.0123 -0.2829 -0.3307 0.1941 0.2058 0.8394 0.2593 0.3625 1.6677 0.3238
Oil Refining/Marketing 9 -0.0159 0.1866 0.6283 0.2447 -0.0744 -0.1300 0.2857 -0.1153 -0.5105 0.3784
Hotels/Resorts 10 0.0259 0.7732 3.1388 0.1350 0.5796 3.2348 0.2466 0.4231 2.9367 0.3681
Restaurants 7 -0.0202 -0.0501 -0.0017 0.1899 -0.1252 -0.5018 0.3136 -0.0738 -0.3704 0.5521
Department/Specialty Retail Stores 8 -0.0060 -0.1601 -0.1885 0.1603 -0.0338 0.0994 0.3261 -0.0367 -0.0017 0.4595
Marine Transportation 14 -0.0003 0.6801 1.3866 0.1603 -0.0446 0.0176 0.2203 0.1103 0.5606 0.2719
EDP Services 11 0.0025 0.4429 1.0640 0.1899 0.0527 0.3354 0.2445 0.0562 0.3661 0.3146
Hospital/Nursing Management 10 -0.0374 0.5054 0.9916 0.1941 -0.0250 0.1177 0.2586 0.0100 0.1850 0.3317
Savings Institutions 6 0.0412 0.5441 2.1445 0.2236 0.3033 1.9788 0.4058 0.1754 1.4319 0.7565
Investment Bankers/Brokers/Service 13 -0.0064 0.2290 0.8349 0.1181 -0.0216 -0.0215 0.1873 0.0205 0.2107 0.2678
Meat/Poultry/Fish 5 -0.0151 -0.2439 -0.6973 0.2743 0.0403 0.3354 0.4512 0.0990 0.8471 1.0000
Beverages (Production/Distribution) 12 0.0239 0.2351 1.0409 0.1013 0.2266 1.5338 0.2250 0.2515 1.9701 0.2991
Homebuilding 16 0.0310 1.6988 3.3707 0.1603 0.9321 3.4975 0.1765 1.0189 4.3494 0.2300
Shoe Manufacturing 5 0.0429 1.4649 1.9803 0.2658 0.4093 1.3059 0.5056 0.3139 1.3117 1.0000
Construction/Ag Equipment/Trucks 6 0.0423 0.5257 1.4138 0.1772 0.5199 2.0191 0.3840 0.4947 2.1520 0.7582
Real Estate 13 0.0009 -0.0317 -0.0225 0.0759 0.1550 1.0338 0.1716 0.1310 1.0304 0.2399
Package Goods/Cosmetics 10 -0.0003 -0.0312 0.0688 0.1477 0.0314 0.2598 0.2396 0.2293 1.3812 0.3537
Oilfield Services/Equipment 8 0.0162 -0.1475 -0.2068 0.1772 -0.3710 -1.2420 0.3026 -0.3021 -1.2370 0.4645
Paper 8 0.0141 0.6179 1.8996 0.2405 0.2799 1.4098 0.3134 0.1059 0.6770 0.4843
Mining & Quarrying of Nonmetallic Minerals (No Fuels) 6 0.0468 -0.0296 0.1624 0.2954 0.0354 0.2772 0.4271 0.0090 0.1427 0.7184
Railroads 10 -0.0005 0.4472 1.7451 0.1477 0.3417 1.6340 0.2414 0.2757 1.6390 0.3544
Home Furnishings 8 0.0309 0.3454 1.1039 0.2025 0.2405 1.0716 0.3136 0.1870 1.0300 0.4658
Apparel 10 -0.0188 -0.5308 -1.3461 0.1603 0.0835 0.4380 0.2529 0.1563 0.7771 0.3381
Office Equipment/Supplies/Services 9 0.0418 0.4468 1.4524 0.1646 -0.0953 -0.3830 0.2719 -0.0194 -0.0141 0.4290
Forest Products 5 0.0311 0.9493 2.0567 0.2785 0.5285 2.1856 0.4836 0.3948 1.9793 1.0000
Auto Manufacturing 9 0.0282 0.5113 1.4388 0.2152 0.1761 0.9958 0.2564 0.3446 2.0070 0.4000
Services-Misc. Amusement & Recreation 6 0.0787 0.8940 2.7186 0.2616 0.5470 3.0031 0.4278 0.3861 2.8966 0.7456
Semiconductors 5 0.0361 1.1884 2.5332 0.2489 0.4766 1.9157 0.4909 0.3035 1.5702 1.0000
Average 10.97 0.0162 0.4170 1.0755 0.1791 0.2574 1.2367 0.2841 0.2257 1.3186 0.4649
Table 8. Performance of baseline model proposed in (Feng et al., 2019) on NYSE
Sector Stocks $\tau$ IRR@1 SR@1 MAP@1 IRR@3 SR@3 MAP@3 IRR@5 SR@5 MAP@5
Computer Manufacturing 5 0.0143 -0.0477 0.0634 0.2658 0.1977 1.0821 0.4838 0.1648 1.0999 1.0000
Transportation Services 5 0.0303 -0.3902 -1.3876 0.2405 -0.3105 -1.6243 0.4913 0.0701 0.4840 1.0000
Medical/Dental Instruments 17 0.0164 1.2784 2.6419 0.0591 0.2390 1.0163 0.1275 0.0349 0.2789 0.1781
Property-Casualty Insurers 19 0.0438 -0.4535 -0.9007 0.1392 -0.2711 -1.1097 0.1636 -0.2591 -1.3919 0.1832
Medical Specialities 5 0.0286 -0.3505 -0.5935 0.2321 -0.3086 -1.5448 $0$0.1143 -0.6432 1.0000
Biotechnology: Biological Products (No Diagnostic Substances) 12 0.0167 -0.5780 -0.9332 0.1224 -0.3050 -1.1125 0.1941 -0.1622 -0.6921 0.2982
Telecommunications Equipment 16 -0.0026 -0.7386 -1.5655 0.1435 -0.4117 -1.4491 0.1535 -0.3187 -1.4648 0.1932
Medical/Nursing Services 6 -0.0149 -0.0497 -0.0130 0.2110 0.1909 0.7797 0.3657 0.1880 0.9182 0.7504
Water Supply 9 0.0138 -0.2709 -0.4976 0.1181 0.2231 1.0461 0.2353 0.2432 1.2927 0.4097
Real Estate Investment Trusts 8 0.0090 -0.3951 -2.2191 0.1139 -0.0881 -0.5507 0.2759 -0.0085 -0.0022 0.4790
Business Services 17 0.0134 0.7775 1.8770 0.0759 0.1112 0.5812 0.1421 0.2183 1.2653 0.2053
Major Pharmaceuticals 29 -0.0131 -0.3425 -0.1044 0.0633 0.1920 0.6252 0.1113 0.2564 0.8491 0.1307
Farming/Seeds/Milling 6 0.0070 -0.2211 -0.6017 0.1435 -0.0610 -0.2920 0.3863 0.0426 0.3586 0.7456
Industrial Specialties 11 0.0141 1.9566 3.0194 0.0844 1.2322 4.2824 0.1960 0.8835 4.1077 0.3188
Air Freight/Delivery Services 6 -0.0120 -0.0647 -0.0398 0.1603 -0.1099 -0.4199 0.3741 -0.0872 -0.3385 0.7380
Electrical Products 13 0.0040 0.2195 0.7759 0.0802 -0.0645 -0.2118 0.1458 -0.0673 -0.3283 0.2409
Television Services 8 0.0048 -0.4989 -1.9251 0.1097 -0.3301 -1.6766 0.2724 -0.2817 -1.6943 0.4904
Investment Bankers/Brokers/Service 18 0.0049 -0.4233 -1.4110 0.1435 -0.4133 -2.8575 0.1620 -0.2782 -2.1985 0.1965
Catalog/Specialty Distribution 7 0.0018 0.0361 0.3542 0.1392 0.4392 1.4817 0.3129 0.3009 1.3728 0.5969
Savings Institutions 20 0.0073 -0.5940 -2.4198 0.0380 -0.2556 -1.3820 0.0956 -0.1932 -1.1429 0.1436
Other Consumer Services 9 0.0184 -0.5178 -1.9858 0.1688 -0.2258 -1.0093 0.2571 -0.0715 -0.3173 0.4181
Auto Parts:O.E.M. 6 0.0048 0.1345 0.5066 0.1730 0.0763 0.4337 0.3884 -0.0166 0.0164 0.7414
Trucking Freight/Courier Services 11 0.0228 0.0093 0.2164 0.1519 0.0257 0.2385 0.2135 0.0142 0.1851 0.3134
Radio And TV Broadcasting And Communications Equipment 10 -0.0251 1.3940 2.4058 0.1139 0.5622 1.8837 0.2053 0.4707 1.9249 0.3386
Biotechnology: Electromedical & Electrotherapeutic Apparatus 5 -0.0219 0.3264 1.1313 0.1646 0.5856 2.5303 0.4655 0.3959 2.0268 1.0000
Home Furnishings 8 0.0183 -0.7652 -2.4810 0.1308 -0.4145 -1.7633 0.2893 -0.2704 -1.3645 0.4916
Major Chemicals 7 0.0006 0.4466 1.4170 0.1308 -0.1249 -0.6651 0.3113 -0.1602 -0.9885 0.5976
Other Specialty Stores 8 -0.0114 -0.6167 -2.5278 0.1392 -0.3897 -2.4056 0.2696 -0.2322 -1.4418 0.4623
Computer Software: Programming, Data Processing 7 -0.0158 0.1778 0.6673 0.2025 0.0463 0.3354 0.3059 0.1507 0.9838 0.5630
Restaurants 20 -0.0096 -0.3113 -1.1495 0.0675 -0.2282 -1.0070 0.1078 -0.1540 -0.7702 0.1484
Automotive Aftermarket 6 -0.0081 -0.1409 -0.4848 0.1646 0.0733 0.4952 0.3854 0.0569 0.4265 0.7266
Investment Managers 5 -0.0412 -0.0107 0.1591 0.1899 0.0263 0.2434 0.4667 0.0409 0.3532 1.0000
Oil Refining/Marketing 6 0.0143 0.6837 1.9976 0.1266 0.2552 1.4019 0.3999 0.0009 0.0929 0.7491
Real Estate 7 0.0154 0.2837 1.1850 0.1181 0.1146 0.7490 0.3291 0.2638 1.7563 0.5990
Apparel 8 -0.0267 -0.1962 -0.4159 0.1435 -0.2422 -1.1042 0.2485 -0.3126 -1.7706 0.4580
Oil & Gas Production 6 -0.0177 0.0560 0.3511 0.1603 -0.1765 -0.4981 0.3704 -0.2671 -1.0112 0.7382
Computer Communications Equipment 5 0.0395 0.3573 1.1938 0.2405 0.1316 0.8054 0.5035 0.0113 0.1607 1.0000
Clothing/Shoe/Accessory Stores 8 0.0252 -0.5559 -1.6409 0.1519 -0.5269 -2.1383 0.2799 -0.3526 -1.4378 0.4914
Commercial Banks 5 0.0580 -0.1046 -0.2805 0.2236 0.0350 0.2759 0.4944 0.1142 0.7075 1.0000
Hospital/Nursing Management 5 0.0092 -0.0562 -0.0759 0.2405 0.1333 0.7221 0.4655 0.1524 0.9918 1.0000
Banks 5 0.0059 -0.2917 -1.6896 0.2025 -0.1515 -0.9863 0.4880 0.1027 0.6914 1.0000
Professional Services 6 0.0188 -0.5588 -2.1915 0.1224 -0.4077 -2.9101 0.3978 0.0101 0.1476 0.7492
Packaged Foods 7 0.0154 1.0455 2.4185 0.1688 0.2836 1.3385 0.3228 -0.0336 -0.1327 0.5925
Metal Fabrications 12 0.0239 1.4819 2.5295 0.1139 0.9541 2.9503 0.1934 0.2094 1.0487 0.2905
Diversified Commercial Services 6 0.0048 -0.5072 -1.6332 0.2321 -0.0854 -0.3199 0.3924 0.2212 1.3537 0.7200
Computer peripheral equipment 6 0.0546 1.0655 2.2080 0.1983 0.1955 0.9741 0.4048 0.1572 0.9538 0.7455
Steel/Iron Ore 7 0.0054 -0.2630 -0.6538 0.1688 -0.0995 -0.3004 0.3380 -0.0112 0.0507 0.5844
Hotels/Resorts 5 -0.0235 -0.1496 -0.4177 0.2152 0.2568 1.2714 0.4735 0.6803 3.0630 1.0000
Biotechnology: In Vitro & In Vivo Diagnostic Substances 5 0.0454 -0.2929 -0.5666 0.2278 0.5973 2.0998 0.5101 0.3823 1.9659 1.0000
Electronic Components 8 -0.0312 0.5785 1.3210 0.1603 0.5248 1.9204 0.2548 0.3378 1.5919 0.4711
Average 9.12 0.0071 0.0310 -0.0873 0.1539 0.0340 0.0445 0.3144 0.0505 0.2678 0.5858
Table 9. Complete result table for LambdaMART on NASDAQ.
Sector Stocks $\tau$ IRR@1 SR@1 MAP@1 IRR@3 SR@3 MAP@3 IRR@5 SR@5 MAP@5
Biotechnology: Laboratory Analytical Instruments 5 0.0000 0.6418 3.0829 0.1899 0.4841 3.4649 0.4970 0.4841 3.4492 1.0000
Diversified Commercial Services 9 0.0128 1.5345 2.2816 0.1983 0.6707 2.3281 0.2647 0.4743 2.2261 0.4146
Other Specialty Stores 11 0.0016 -0.4049 -0.7317 0.1435 -0.2758 -0.8344 0.2161 -0.1138 -0.3821 0.3255
Investment Managers 18 0.0280 -0.0565 -0.0769 0.0970 0.2047 1.1636 0.1484 0.3683 2.2241 0.1888
Electrical Products 13 0.0080 -0.0764 -0.2386 0.1055 -0.0166 -0.0226 0.1578 0.0685 0.5614 0.2450
Major Pharmaceuticals 18 0.0318 0.7983 1.4109 0.1308 0.1837 0.8006 0.1835 0.1154 0.6883 0.2122
Other Pharmaceuticals 5 -0.0143 -0.0489 -0.0297 0.2025 -0.0429 -0.1446 0.4712 0.1293 0.8351 1.0000
Automotive Aftermarket 10 0.0268 0.3631 1.0532 0.0886 -0.0249 0.0486 0.2075 -0.0197 0.0353 0.3614
Precious Metals 11 0.0147 -0.2612 -0.9794 0.1139 -0.2128 -1.0622 0.2063 0.0052 0.1263 0.3314
Publishing 5 0.0479 -0.0885 -0.1322 0.2447 -0.0158 0.0293 0.5084 -0.1535 -0.7639 1.0000
Military/Government/Technical 11 -0.0206 0.3135 1.2221 0.1181 0.2572 1.3468 0.1880 0.2948 1.7646 0.2961
Business Services 22 0.0118 -0.2033 -1.0205 0.0844 -0.3400 -2.8763 0.1045 -0.2780 -2.4584 0.1312
Packaged Foods 16 0.0146 -0.0450 -0.0991 0.0675 -0.1055 -0.6338 0.1153 -0.0492 -0.3276 0.1915
Life Insurance 17 0.0013 0.4844 2.2569 0.0338 0.9186 5.5226 0.1235 0.7000 4.8547 0.1734
Power Generation 15 0.0001 0.4149 1.4183 0.1055 0.1330 1.0341 0.1486 0.0962 0.9612 0.2192
Clothing/Shoe/Accessory Stores 15 -0.0152 -0.3638 -1.1112 0.0802 -0.5180 -2.2291 0.1378 -0.4522 -1.8764 0.1949
Medical Specialities 11 0.0163 0.1157 0.9983 0.0549 0.0511 0.5236 0.2002 0.0434 0.4638 0.3105
Property-Casualty Insurers 30 -0.0150 -0.3339 -1.8750 0.0253 -0.2293 -1.7631 0.0539 -0.1881 -1.6600 0.0851
Accident &Health Insurance 5 0.0193 0.0647 0.3770 0.2658 0.1197 0.8376 0.4862 0.2289 1.5754 1.0000
Farming/Seeds/Milling 7 0.0210 -0.2092 -1.0287 0.1730 -0.0773 -0.5290 0.3408 -0.0765 -0.6195 0.5825
Engineering & Construction 5 -0.0286 -0.6108 -1.9237 0.2236 -0.3614 -1.7340 0.4709 -0.0749 -0.2026 1.0000
Aerospace 9 0.0138 -0.0585 -0.0292 0.1435 0.2664 1.3216 0.2480 0.2080 1.3286 0.4100
Industrial Specialties 9 -0.0093 0.2417 0.7658 0.1308 0.3355 1.7741 0.2419 0.4916 2.9676 0.3970
Specialty Insurers 6 0.0132 0.0248 0.2510 0.1772 0.1812 1.7888 0.3905 0.1476 1.5814 0.7401
Major Chemicals 29 -0.0014 0.9222 2.1943 0.0633 0.2557 1.3246 0.0759 0.2242 1.4245 0.1104
Integrated oil Companies 19 0.0107 0.1442 0.6068 0.1181 0.2332 1.1554 0.1397 0.2060 1.1236 0.1846
Air Freight/Delivery Services 8 0.0123 0.5049 1.8273 0.1603 -0.0349 -0.0056 0.3019 -0.0296 -0.0516 0.4711
Auto Parts:O.E.M. 18 -0.0036 0.3236 1.1766 0.0633 0.3508 1.5749 0.1041 0.4035 1.9846 0.1615
Metal Fabrications 23 0.0044 0.2061 0.7621 0.0844 0.4328 1.6071 0.1085 0.2989 1.3637 0.1329
Professional Services 11 -0.0002 0.0308 0.2546 0.1392 0.1195 0.7118 0.2135 0.1413 0.9830 0.3149
Consumer Electronics/Appliances 6 0.0244 0.4636 1.7942 0.1899 0.3072 2.2109 0.3929 0.1913 1.7115 0.7528
Fluid Controls 7 0.0094 -0.2719 -0.7571 0.1435 0.0033 0.1266 0.3298 0.0295 0.2601 0.5846
Other Consumer Services 17 0.0155 -0.1323 0.0756 0.1097 -0.2485 -0.9130 0.1418 -$0$0.0941 0.1920
Steel/Iron Ore 9 0.0121 0.1835 0.5921 0.1266 0.2446 0.8770 0.2471 0.2832 1.1043 0.4234
Oil/Gas Transmission 6 -0.0165 0.0695 0.5316 0.1477 0.1606 1.4076 0.3868 0.1223 1.2065 0.7418
Plastic Products 6 -0.0003 -0.1394 -0.2104 0.1814 0.0448 0.3587 0.3715 0.2178 1.5464 0.7333
Agricultural Chemicals 7 0.0262 -0.1353 -0.3462 0.1561 0.1867 0.9687 0.3371 0.0902 0.6374 0.5874
Containers/Packaging 9 0.0100 0.4112 2.4134 0.1013 0.2591 2.1785 0.2278 0.2204 1.9453 0.4089
Water Supply 7 0.0302 -0.3861 -1.7638 0.1814 -0.1310 -0.6964 0.3469 0.0101 0.1464 0.5939
Finance: Consumer Services 20 0.0096 0.9553 2.8125 0.0422 0.5232 2.5844 0.1184 0.2217 1.4814 0.1665
Commercial Banks 21 0.0075 0.1036 0.5447 0.0506 0.4271 2.3475 0.1055 0.4649 3.0711 0.1418
Medical/Dental Instruments 12 0.0127 -0.2963 -1.7528 0.1139 -0.2316 -2.0540 0.1707 -0.2241 -2.1604 0.2663
Computer Software: Prepackaged Software 11 0.0120 -0.4054 -1.0410 0.0802 -0.0426 -0.0959 0.1878 0.0912 0.6130 0.3252
Oil Refining/Marketing 9 0.0049 -0.4859 -1.2104 0.0759 -0.2478 -1.1750 0.2351 -0.2189 -1.3731 0.4055
Hotels/Resorts 10 -0.0106 -0.1800 -0.6967 0.1224 -0.0396 -0.1907 0.2023 0.1254 0.9986 0.3454
Restaurants 7 -0.0126 -0.4359 -2.0032 0.1519 -0.0917 -0.5576 0.3139 -0.0712 -0.4825 0.5702
Department/Specialty Retail Stores 8 0.0051 0.0043 0.2435 0.1857 -0.1060 -0.1541 0.3040 -0.1186 -0.3509 0.4650
Marine Transportation 14 0.0335 -0.2724 -0.6564 0.0970 -0.2722 -1.0105 0.1756 -0.1424 -0.6389 0.2513
EDP Services 11 -0.0057 -0.1577 -0.4264 0.1055 0.1571 0.9432 0.1845 0.2683 1.5957 0.2959
Hospital/Nursing Management 10 0.0193 4.2295 3.5191 0.1097 0.8279 2.2462 0.2389 0.5408 1.9074 0.3636
Savings Institutions 6 0.0205 -0.0950 -0.5292 0.1730 0.0393 0.3589 0.4001 0.0866 0.7941 0.7442
Investment Bankers/Brokers/Service 13 -0.0240 0.2734 1.0000 0.0802 0.0255 0.2335 0.1606 -0.0393 -0.1499 0.2442
Meat/Poultry/Fish 5 -0.0135 -0.0616 -0.1277 0.2363 0.1429 1.0507 0.4669 0.1074 0.9181 1.0000
Beverages (Production/Distribution) 12 0.0299 -0.1384 -0.6496 0.0802 -0.0516 -0.3698 0.1929 -0.0390 -0.3551 0.2900
Homebuilding 16 0.0151 -0.2566 -1.3115 0.0759 0.0165 0.1893 0.1428 0.0038 0.1021 0.2014
Shoe Manufacturing 5 0.0361 -0.3162 -0.5245 0.2236 -0.1060 -0.2621 0.5164 0.2378 1.0544 1.0000
Construction/Ag Equipment/Trucks 6 -0.0373 0.1895 0.7694 0.1857 0.3942 1.5671 0.3788 0.4011 1.8141 0.7335
Real Estate 13 -0.0286 -0.6280 -3.5603 0.0802 -0.3128 -2.2656 0.1423 -0.1233 -0.9755 0.2367
Package Goods/Cosmetics 10 -0.0092 -0.0532 -0.0795 0.1013 0.1455 1.1054 0.2028 0.1955 1.6992 0.3396
Oilfield Services/Equipment 8 -0.0315 -0.2869 -0.8246 0.1181 -0.4142 -1.9334 0.2386 -0.3447 -1.6829 0.4587
Paper 8 0.0141 -0.2248 -0.8357 0.1266 -0.1451 -0.7288 0.2860 0.0000 0.0900 0.4747
Mining & Quarrying of Nonmetallic Minerals (No Fuels) 6 0.0389 2.0392 4.2855 0.1730 0.5500 2.5453 0.3837 0.1273 0.8274 0.7666
Railroads 10 -0.0085 0.1485 0.6666 0.1055 0.1695 1.0637 0.1922 0.1236 0.9389 0.3336
Home Furnishings 8 0.0165 -0.3182 -0.7994 0.1477 -0.0887 -0.3797 0.2792 -0.0377 -0.1389 0.4869
Apparel 10 -0.0281 -0.4196 -0.8942 0.1350 -0.1056 -0.3244 0.2353 0.0245 0.2272 0.3297
Office Equipment/Supplies/Services 9 0.0301 0.1822 0.7178 0.1392 -0.0653 -0.1916 0.2836 -0.0758 -0.3325 0.4270
Forest Products 5 -0.0185 0.0951 0.5010 0.1983 0.2182 1.1585 0.4662 0.3766 1.8965 1.0000
Auto Manufacturing 9 -0.0187 0.1151 0.5711 0.1055 0.0523 0.4158 0.2260 0.0694 0.5413 0.3996
Services-Misc. Amusement & Recreation 6 0.0182 -0.2761 -1.3759 0.1899 0.0169 0.1929 0.4046 0.1916 1.5817 0.7536
Semiconductors 5 -0.0118 -0.4208 -1.7579 0.1730 -0.0284 -0.0098 0.4759 0.4026 2.0180 1.0000
Average 10.97 0.0054 0.1005 0.1367 0.1307 0.0732 0.4192 0.2593 0.1063 0.6882 0.4574
Table 10. Complete result table for LambdaMART on NYSE.
Sector Stocks $\tau$ IRR@1 SR@1 MAP@1 IRR@3 SR@3 MAP@3 IRR@5 SR@5 MAP@5
Computer Manufacturing 5 -0.0244 -0.0458 0.0585 0.2110 0.2554 1.1537 0.4627 0.2167 1.3724 1.0000
Transportation Services 5 -0.0193 -0.3402 -1.1114 0.1941 0.2787 1.2932 0.4557 0.1263 0.7757 1.0000
Medical/Dental Instruments 17 0.0492 0.1590 0.6044 0.1055 0.6454 2.2254 0.1592 0.6764 2.7925 0.2023
Property-Casualty Insurers 19 0.0136 0.1661 0.6221 0.0928 -0.0236 0.0176 0.1306 -0.1705 -0.7351 0.1768
Medical Specialities 5 0.0336 -0.0239 0.1681 0.2405 0.2459 1.1755 0.4768 -0.0402 -0.1500 1.0000
Biotechnology: Biological Products (No Diagnostic Substances) 12 0.0395 1.4993 2.0935 0.1519 1.4277 3.3088 0.2429 0.5337 1.8510 0.3092
Telecommunications Equipment 16 -0.0179 -0.7585 -1.6430 0.1688 -0.2981 -0.7723 0.1779 -0.1275 -0.3238 0.2197
Medical/Nursing Services 6 -0.0311 0.1104 0.4698 0.1561 0.3651 1.4513 0.3659 0.3374 1.5209 0.7409
Water Supply 9 0.0574 0.4483 1.3802 0.2068 0.5629 2.3886 0.2843 0.3641 1.8486 0.4219
Real Estate Investment Trusts 8 -0.0132 0.0069 0.1270 0.0844 -0.0191 -0.0712 0.2595 -0.0638 -0.4460 0.4922
Business Services 17 0.0253 0.4529 1.0647 0.1392 0.5424 1.9924 0.1655 0.6901 2.8825 0.2158
Major Pharmaceuticals 29 0.0079 -0.1586 -0.0520 0.0506 0.3816 0.8635 0.1174 0.5645 1.3108 0.1410
Farming/Seeds/Milling 6 0.0216 0.3578 1.1610 0.1899 0.1054 0.6186 0.3685 0.1074 0.7423 0.7387
Industrial Specialties 11 0.0355 2.0407 2.9100 0.0886 1.1427 3.4837 0.2028 0.8774 3.5009 0.3327
Air Freight/Delivery Services 6 0.0008 0.0305 0.2780 0.1983 -0.0337 -0.0015 0.3706 0.0795 0.4713 0.7377
Electrical Products 13 0.0238 0.7368 1.5589 0.0675 0.5567 1.8610 0.1782 0.7503 2.7229 0.2683
Television Services 8 -0.0102 -0.1655 -0.3784 0.1266 -0.1655 -0.5383 0.2979 -0.1354 -0.5403 0.4825
Investment Bankers/Brokers/Service 18 0.0182 0.9022 2.0099 0.1688 0.1769 1.1874 0.2114 0.1941 1.2786 0.2410
Catalog/Specialty Distribution 7 0.0354 0.2672 0.7606 0.1435 0.7167 1.8853 0.3401 0.5044 1.9389 0.5866
Savings Institutions 20 0.0530 0.5454 1.5980 0.1224 0.2679 1.2659 0.1615 0.2020 1.1164 0.1943
Other Consumer Services 9 -0.0166 -0.0931 -0.2221 0.1055 0.1196 0.7069 0.2257 0.1118 0.7043 0.3956
Auto Parts:O.E.M. 6 0.0345 -0.1027 -0.3203 0.1055 -0.0189 0.0618 0.3987 -0.0246 -0.0228 0.7652
Trucking Freight/Courier Services 11 0.0244 0.4114 0.9919 0.1899 0.4502 1.3182 0.2658 0.4701 1.4809 0.3332
Radio And Television Broadcasting And Communications Equipment 10 0.0109 -0.0742 -0.1502 0.0633 0.1295 0.7907 0.1845 -0.0134 0.0003 0.3628
Biotechnology: Electromedical & Electrotherapeutic Apparatus 5 0.0496 1.1587 2.2743 0.2532 0.8958 3.0769 0.5075 0.3760 1.9405 1.0000
Home Furnishings 8 0.0285 -0.2161 -0.4893 0.1435 -0.0517 -0.0146 0.2785 0.0614 0.3821 0.4723
Major Chemicals 7 -0.0138 -0.2516 -0.6296 0.1772 0.1080 0.5532 0.3317 0.0866 0.5661 0.5419
Other Specialty Stores 8 -0.0180 -0.3415 -0.4555 0.2110 -0.2925 -0.7834 0.2745 -0.0670 -0.1054 0.4397
Computer Software: Programming, Data Processing 7 -0.0158 -0.1693 -0.2412 0.2194 0.1802 0.8314 0.3338 0.3037 1.6221 0.5186
Restaurants 20 -0.0124 -0.3350 -0.8094 0.0380 0.1699 0.7861 0.1470 0.1158 0.6462 0.1718
Automotive Aftermarket 6 0.0675 0.9524 2.4494 0.2489 0.3451 1.5889 0.4456 0.2365 1.3197 0.7386
Investment Managers 5 -0.0479 -0.0173 -0.0278 0.1308 0.0694 0.4809 0.4447 0.0976 0.7306 1.0000
Oil Refining/Marketing 6 -0.0177 -0.2807 -1.0433 0.2785 0.0871 0.5231 0.3664 0.1192 0.7835 0.6866
Real Estate 7 0.0346 0.0248 0.2712 0.1435 0.3611 1.8716 0.3226 0.3814 2.2677 0.6013
Apparel 8 -0.0126 -0.1179 -0.4358 0.1055 -0.0198 0.0517 0.2583 0.1292 0.7542 0.4581
Oil & Gas Production 6 -0.0317 -0.3547 -0.9233 0.1688 -0.3303 -1.1890 0.3507 -0.3033 -1.1135 0.7115
Computer Communications Equipment 5 0.0084 1.4289 2.4003 0.2405 0.3316 1.4066 0.4632 0.0383 0.3025 1.0000
Clothing/Shoe/Accessory Stores 8 -0.0123 -0.3926 -0.7858 0.1435 -0.1718 -0.3006 0.2778 -0.0946 -0.1198 0.4774
Commercial Banks 5 0.0008 0.2494 0.8646 0.2236 0.2425 1.0891 0.4749 0.1656 0.9570 1.0000
Hospital/Nursing Management 5 0.0168 0.4520 1.3993 0.2194 0.5566 2.6969 0.5056 0.2216 1.3716 1.0000
Banks 5 0.0975 0.4188 1.6950 0.2194 0.3821 1.9376 0.5333 0.0939 0.6403 1.0000
Professional Services 6 0.0367 -0.1459 -0.4075 0.1646 0.2578 1.3598 0.3896 0.2896 1.7259 0.7525
Packaged Foods 7 0.0262 -0.0358 0.0393 0.1561 0.0106 0.1580 0.3479 0.0664 0.4664 0.5737
Metal Fabrications 12 0.0107 0.8319 1.7550 0.1266 0.3500 1.3203 0.2131 0.4291 1.7411 0.2848
Diversified Commercial Services 6 0.0451 0.1541 0.5682 0.2743 0.5603 2.3008 0.4030 0.3485 2.0433 0.7311
Computer peripheral equipment 6 0.0076 0.2860 0.8418 0.2152 0.4543 1.7633 0.3826 0.3222 1.6209 0.7227
Steel/Iron Ore 7 -0.0190 -0.2737 -0.9245 0.1097 -0.1345 -0.4884 0.2989 -0.0798 -0.2852 0.5987
Hotels/Resorts 5 -0.0109 0.5641 2.9180 0.1646 0.9370 3.6587 0.4780 0.8753 3.8558 1.0000
Biotechnology: In Vitro & In Vivo Diagnostic Substances 5 -0.0151 0.9486 2.0263 0.2110 0.4172 1.5755 0.4665 0.4424 2.1666 1.0000
Electronic Components 8 -0.0069 0.3731 1.0121 0.1392 0.7566 2.3061 0.2789 0.6594 2.3433 0.4611
Average 9.12 0.0110 0.2257 0.5464 0.1620 0.2857 1.1245 0.3216 0.2309 1.0943 0.5860
Table 11. Complete result table for Rankformer on NASDAQ.
Sector Stocks $\tau$ IRR@1 SR@1 MAP@1 IRR@3 SR@3 MAP@3 IRR@5 SR@5 MAP@5
Biotechnology: Laboratory Analytical Instruments 5 -0.0025 0.6902 2.1537 0.1983 0.5843 3.4616 0.4698 0.4553 3.1498 1.0000
Diversified Commercial Services 9 0.0367 1.4739 1.9763 0.2405 1.0969 2.9188 0.3043 0.8108 3.0485 0.4004
Other Specialty Stores 11 0.0059 -0.1341 -0.0301 0.1519 0.3229 1.1175 0.2370 0.0706 0.4060 0.3223
Investment Managers 18 0.0376 0.7707 2.4018 0.1139 0.6134 3.1095 0.1526 0.5251 3.1464 0.2086
Electrical Products 13 0.0253 0.5033 1.4703 0.1097 0.1957 1.1061 0.1695 0.3724 2.1064 0.2651
Major Pharmaceuticals 18 0.0117 -0.1838 -0.0680 0.1772 -0.1492 -0.3752 0.2199 0.0132 0.1763 0.2236
Other Pharmaceuticals 5 -0.0034 0.3746 1.4293 0.2405 0.1998 1.1667 0.4583 0.1719 1.0601 1.0000
Automotive Aftermarket 10 0.0421 0.1357 0.6239 0.0591 0.2499 0.9958 0.2112 0.2311 0.9409 0.3905
Precious Metals 11 -0.0154 -0.2354 -0.4765 0.1646 -0.1428 -0.5177 0.1969 -0.0027 0.0953 0.3068
Publishing 5 -0.0160 0.0131 0.1672 0.1013 -0.0435 -0.1085 0.4831 -0.1218 -0.5510 1.0000
Military/Government/Technical 11 0.0274 0.3260 1.0909 0.1181 0.2801 1.3298 0.1887 0.2861 1.5777 0.3309
Business Services 22 0.0470 0.5503 1.6952 0.0675 0.6726 3.2061 0.1472 0.6842 4.0307 0.1781
Packaged Foods 16 0.0202 0.0653 0.4200 0.1097 0.3028 1.2772 0.1920 0.2975 1.5563 0.2549
Life Insurance 17 0.0321 0.7816 1.8057 0.1814 0.3132 1.4290 0.1920 0.3749 1.9991 0.2148
Power Generation 15 0.0293 0.0506 0.3164 0.1097 0.1053 0.7563 0.1774 0.1561 1.3379 0.2333
Clothing/Shoe/Accessory Stores 15 0.0225 0.6554 1.1923 0.0759 0.1537 0.5621 0.1535 0.1204 0.5030 0.2334
Medical Specialities 11 0.0630 0.1184 0.5023 0.2236 0.5825 2.4361 0.2487 0.5778 3.0334 0.3436
Property-Casualty Insurers 30 0.0222 0.3422 0.8457 0.1561 0.0831 0.4595 0.1430 0.0884 0.5538 0.1626
Accident &Health Insurance 5 0.0160 0.2131 1.1893 0.1899 0.2929 1.9963 0.4965 0.2645 1.7726 1.0000
Farming/Seeds/Milling 7 0.0458 0.7947 2.8643 0.2110 0.1444 0.9477 0.3516 0.0944 0.7830 0.5795
Engineering & Construction 5 -0.0177 -0.0470 -0.0562 0.1857 0.0668 0.4253 0.4752 -0.0339 -0.0130 1.0000
Aerospace 9 0.0070 0.0075 0.2826 0.2110 0.2110 0.9805 0.2700 0.3692 1.9613 0.3847
Industrial Specialties 9 0.0233 0.5945 1.2360 0.1814 0.4902 1.9122 0.2867 0.4382 2.3662 0.3906
Specialty Insurers 6 0.0266 0.2417 1.9522 0.1772 0.2524 2.3140 0.4182 0.1542 1.5739 0.7641
Major Chemicals 29 0.0385 0.7861 2.3038 0.0506 0.9033 2.5378 0.1083 0.9413 3.3278 0.1439
Integrated oil Companies 19 0.0030 -0.1368 -0.1112 0.1181 0.3038 1.2159 0.1861 0.2780 1.2721 0.2256
Air Freight/Delivery Services 8 -0.0264 -0.1159 -0.4758 0.1224 0.0035 0.1200 0.2447 0.0342 0.2842 0.5034
Auto Parts:O.E.M. 18 0.0093 -0.0773 -0.1814 0.0169 0.1225 0.6453 0.1027 0.2304 1.2174 0.1590
Metal Fabrications 23 0.0315 0.4166 1.2780 0.0464 0.2232 1.1539 0.0748 0.3343 1.8060 0.1225
Professional Services 11 0.0424 0.3269 1.2574 0.0928 0.7564 3.3393 0.2011 0.4197 2.2984 0.3334
Consumer Electronics/Appliances 6 0.0143 -0.0331 -0.0448 0.2236 0.1965 1.3905 0.3849 0.1815 1.6288 0.7276
Fluid Controls 7 0.0046 0.2013 0.8376 0.1308 0.0164 0.1977 0.3425 0.1301 0.7341 0.5530
Other Consumer Services 17 0.0112 -0.1938 -0.3318 0.1055 0.0592 0.3613 0.1786 0.0714 0.4484 0.2146
Steel/Iron Ore 9 0.0063 0.4187 0.9607 0.1308 0.4679 1.4640 0.2604 0.3296 1.2657 0.4055
Oil/Gas Transmission 6 0.0429 0.0450 0.3377 0.1772 0.1682 1.3444 0.4004 0.1402 1.2989 0.7125
Plastic Products 6 0.0445 0.5481 1.7984 0.1983 0.2097 1.1128 0.4119 0.2402 1.6555 0.7280
Agricultural Chemicals 7 -0.0274 -0.1482 -0.5050 0.1139 0.0579 0.3661 0.3099 0.0235 0.2252 0.5530
Containers/Packaging 9 0.0560 0.3614 1.6604 0.1857 0.4993 3.0729 0.2811 0.4182 3.0827 0.4291
Water Supply 7 0.0434 0.3842 1.2216 0.2321 0.4928 2.1438 0.3462 0.5040 2.6645 0.5847
Finance: Consumer Services 20 0.0222 -0.0371 0.0167 0.0928 0.3712 1.6133 0.1367 0.3801 1.8747 0.1840
Commercial Banks 21 0.0191 0.1958 0.9104 0.0759 0.2458 1.3927 0.1327 0.2432 1.5127 0.1786
Medical/Dental Instruments 12 0.0359 0.9956 3.0300 0.1139 0.7530 3.7000 0.2131 0.6578 3.9586 0.3012
Computer Software: Prepackaged Software 11 0.0051 0.3920 1.8192 0.0802 0.3935 2.3690 0.1770 0.3880 2.4915 0.2962
Oil Refining/Marketing 9 -0.0131 -0.0392 0.2148 0.2194 -0.0582 -0.0842 0.2806 -0.1034 -0.4494 0.3798
Hotels/Resorts 10 0.0302 0.5783 2.1517 0.1603 0.4510 2.6568 0.2363 0.4765 3.1656 0.3614
Restaurants 7 -0.0250 -0.0055 0.1315 0.1941 -0.1752 -0.7641 0.3181 -0.0880 -0.4705 0.5514
Department/Specialty Retail Stores 8 -0.0093 -0.0487 -0.6587 0.1181 0.0553 0.4820 0.2581 0.1323 0.9452 0.5083
Marine Transportation 14 -0.0028 0.3115 0.8296 0.1561 -0.0475 0.0074 0.2211 0.1118 0.5671 0.2705
EDP Services 11 0.0016 0.3955 0.9938 0.1857 0.2148 0.8664 0.2482 0.0830 0.4807 0.3157
Hospital/Nursing Management 10 0.0094 0.3321 1.0789 0.1139 0.0799 0.4927 0.2016 0.0709 0.4200 0.3503
Savings Institutions 6 0.0233 0.5687 2.2148 0.1983 0.2700 1.7638 0.4015 0.1603 1.3813 0.7545
Investment Bankers/Brokers/Service 13 -0.0005 0.2942 1.0115 0.1181 0.0082 0.1411 0.1854 0.0205 0.2107 0.2678
Meat/Poultry/Fish 5 -0.0143 -0.2439 -0.6973 0.2743 0.0873 0.6218 0.4543 0.0990 0.8471 1.0000
Beverages (Production/Distribution) 12 0.0192 0.1427 0.6925 0.1055 0.2875 1.9098 0.2288 0.2508 1.9447 0.2990
Homebuilding 16 0.0415 1.1992 2.3832 0.1139 0.9380 3.1958 0.1707 1.1283 4.2944 0.2342
Shoe Manufacturing 5 0.0387 0.8778 1.4294 0.2110 0.5701 1.6533 0.5089 0.3139 1.3117 1.0000
Construction/Ag Equipment/Trucks 6 0.0457 0.7575 1.8814 0.1730 0.4618 1.8972 0.3821 0.5135 2.2277 0.7643
Real Estate 13 0.0035 -0.0317 -0.0225 0.0759 0.1757 1.1769 0.1742 0.1773 1.3251 0.2446
Package Goods/Cosmetics 10 0.0098 -0.0312 0.0688 0.1477 0.0474 0.3356 0.2410 0.2505 1.5186 0.3546
Oilfield Services/Equipment 8 -0.0069 -0.4508 -0.9192 0.2321 -0.1833 -0.5156 0.2717 -0.2205 -0.8215 0.4629
Paper 8 0.0150 0.6198 1.8548 0.2152 0.3067 1.4751 0.3050 0.0984 0.6318 0.4842
Mining & Quarrying of Nonmetallic Minerals (No Fuels) 6 0.0569 -0.0299 0.1618 0.2954 0.0305 0.2549 0.4313 -0.0001 0.0885 0.7188
Railroads 10 -0.0109 0.4033 1.3568 0.1266 0.2680 1.2789 0.2410 0.2466 1.4833 0.3504
Home Furnishings 8 0.0336 0.2567 0.8786 0.1899 0.2967 1.3483 0.3080 0.1965 1.0645 0.4663
Apparel 10 0.0079 0.2067 0.8641 0.1055 0.1946 0.9773 0.2025 0.3079 1.4040 0.3496
Office Equipment/Supplies/Services 9 -0.0390 -0.3344 -1.2649 0.0844 -0.3280 -1.6894 0.2274 -0.2768 -1.5305 0.3900
Forest Products 5 0.0664 0.9244 2.0120 0.2532 0.7321 2.7681 0.5052 0.3948 1.9793 1.0000
Auto Manufacturing 9 0.0252 0.0875 0.4246 0.2025 0.1577 0.9115 0.2520 0.3126 1.8921 0.3949
Services-Misc. Amusement & Recreation 6 0.0602 0.5462 2.0205 0.1857 0.4993 2.8017 0.4308 0.3038 2.2588 0.7432
Semiconductors 5 0.0378 0.7516 1.9347 0.2236 0.4438 1.8223 0.4859 0.3035 1.5702 1.0000
Average 10.97 0.0181 0.2924 0.9113 0.1535 0.2701 1.2890 0.2758 0.2515 1.4200 0.465
Table 12. Complete result table for Rankformer on NYSE.
Sector Stocks $\tau$ IRR@1 SR@1 MAP@1 IRR@3 SR@3 MAP@3 IRR@5 SR@5 MAP@5
Computer Manufacturing 5 0.0345 0.1128 0.6215 0.1983 0.3465 2.3286 0.5059 0.2167 1.3724 1.0000
Transportation Services 5 0.0008 0.4351 1.2510 0.2489 0.2339 1.1572 0.4838 0.1263 0.7757 1.0000
Medical/Dental Instruments 17 0.0090 0.5775 1.9145 0.0802 0.4503 2.3065 0.1254 0.4129 2.3442 0.1809
Property-Casualty Insurers 19 0.0194 -0.0452 0.1845 0.1013 0.0247 0.2305 0.1751 0.0401 0.2988 0.1967
Medical Specialities 5 0.0303 -0.0239 0.1681 0.2405 0.0473 0.3352 0.4930 -0.0402 -0.1500 1.0000
Biotechnology: Biological Products (No Diagnostic Substances) 12 0.0186 2.4685 2.9704 0.1688 1.0795 3.2313 0.2117 0.4928 1.8969 0.3013
Telecommunications Equipment 16 0.0312 0.2872 0.9428 0.0675 0.0301 0.2521 0.1292 0.1160 0.6588 0.2023
Medical/Nursing Services 6 0.0345 0.6698 1.5659 0.1688 0.7648 2.4595 0.3915 0.3438 1.4557 0.7572
Water Supply 9 0.0392 0.4756 1.4303 0.2110 0.4821 2.1976 0.2785 0.3584 1.8276 0.4219
Real Estate Investment Trusts 8 0.0063 0.1655 1.0105 0.1392 0.1347 1.1262 0.2818 0.0755 0.6959 0.4834
Business Services 17 0.0268 1.6051 2.4223 0.1603 0.9827 3.3542 0.1798 0.7203 3.4705 0.2180
Major Pharmaceuticals 29 0.0147 1.7511 2.5307 0.0380 1.1160 2.4861 0.0649 0.7787 2.4209 0.0850
Farming/Seeds/Milling 6 0.0479 0.2494 0.8739 0.2152 0.2853 1.4940 0.4044 0.0968 0.6758 0.7522
Industrial Specialties 11 0.0219 1.4803 2.9657 0.0844 0.9133 2.7340 0.2220 0.7123 2.9349 0.3298
Air Freight/Delivery Services 6 0.0165 0.5981 1.5902 0.1772 0.1919 0.8855 0.4121 0.0617 0.3960 0.7350
Electrical Products 13 0.0139 0.7310 2.2919 0.1055 0.3622 1.6007 0.1772 0.4157 1.8738 0.2609
Television Services 8 0.0318 -0.0302 0.0737 0.1181 0.0519 0.3379 0.3061 -0.0576 -0.1870 0.5070
Investment Bankers/Brokers/Service 18 0.0366 0.0700 0.3793 0.1055 0.4111 2.0910 0.1343 0.5326 3.0702 0.1857
Catalog/Specialty Distribution 7 0.0474 1.0474 1.6694 0.1730 0.9757 2.4631 0.3209 0.6368 2.3162 0.6066
Savings Institutions 20 0.0341 0.4459 1.3907 0.1266 0.3911 1.7948 0.1472 0.2374 1.2924 0.1919
Other Consumer Services 9 0.0327 -0.0204 0.1423 0.1350 0.3638 1.6126 0.2714 0.1654 0.8493 0.4181
Auto Parts:O.E.M. 6 0.0205 0.1364 0.5079 0.2278 0.0853 0.4936 0.4083 -0.0303 -0.0533 0.7472
Trucking Freight/Courier Services 11 0.0236 1.2633 1.9697 0.1941 0.8822 2.3363 0.2698 0.4234 1.4514 0.3321
Radio And TV Broadcasting And Communications Equipment 10 0.0020 0.6117 1.1721 0.1097 0.5461 1.8837 0.2218 0.2672 1.3208 0.3429
Biotechnology: Electromedical & Electrotherapeutic Apparatus 5 0.0336 0.3490 1.3161 0.2152 0.6316 2.6888 0.4892 0.3760 1.9405 1.0000
Home Furnishings 8 0.0255 -0.5294 -0.9480 0.1772 -0.0071 0.1231 0.2902 0.0675 0.4162 0.4921
Major Chemicals 7 0.0386 0.0550 0.3581 0.1266 0.2015 1.0694 0.3293 0.1306 0.8491 0.6094
Other Specialty Stores 8 0.0666 0.6364 2.0735 0.1266 0.9138 3.1965 0.3033 0.1963 1.0324 0.5221
Computer Software: Programming, Data Processing 7 0.0470 0.7064 2.3288 0.1814 0.6583 3.4336 0.3305 0.3598 2.1732 0.6190
Restaurants 20 0.0077 0.1668 0.6579 0.0506 0.1775 0.9692 0.0942 0.1184 0.7246 0.1443
Automotive Aftermarket 6 0.0328 0.3689 1.2364 0.2025 0.3357 1.7185 0.3999 0.2355 1.3734 0.7544
Investment Managers 5 0.0471 -0.0055 0.1115 0.2110 0.0272 0.2449 0.5087 0.0976 0.7306 1.0000
Oil Refining/Marketing 6 0.0317 0.5765 1.6158 0.2532 0.3921 1.8925 0.4154 0.1472 0.9318 0.7284
Real Estate 7 0.0454 0.0186 0.2540 0.1350 0.4378 2.0464 0.3376 0.3771 2.2129 0.5996
Apparel 8 0.0366 0.9977 1.9552 0.1688 0.4962 1.6474 0.3054 0.3216 1.4747 0.4992
Oil & Gas Production 6 0.0160 -0.3752 -1.0564 0.1519 -0.0738 -0.0969 0.3826 -0.2398 -0.8349 0.7308
Computer Communications Equipment 5 0.0168 0.8553 2.3743 0.2785 0.2458 1.1800 0.4777 0.0383 0.3025 1.0000
Clothing/Shoe/Accessory Stores 8 0.0582 0.0546 0.3739 0.1477 0.1752 0.6443 0.3064 0.1487 0.6079 0.5038
Commercial Banks 5 0.0361 0.0714 0.4419 0.1603 0.4664 2.2140 0.5415 0.1656 0.9570 1.0000
Hospital/Nursing Management 5 0.0076 0.4818 1.7654 0.2321 0.4274 2.4203 0.4794 0.2216 1.3716 1.0000
Banks 5 0.0227 0.2333 1.0907 0.2194 0.1565 0.9575 0.5002 0.0939 0.6403 1.0000
Professional Services 6 0.0625 0.6336 2.1653 0.2194 0.3393 1.7894 0.4255 0.2238 1.3882 0.7533
Packaged Foods 7 0.0262 -0.0920 -0.1460 0.1730 -0.0812 -0.3737 0.3474 -0.0810 -0.4288 0.5873
Metal Fabrications 12 0.0157 0.1418 0.6525 0.0464 0.4426 1.7620 0.1650 0.1618 0.8427 0.2879
Diversified Commercial Services 6 0.0501 0.6682 2.4055 0.1899 0.5824 2.4056 0.4144 0.1631 1.0603 0.7604
Computer peripheral equipment 6 0.0199 1.3021 2.2781 0.1857 0.7407 2.5088 0.3971 0.2258 1.2091 0.7362
Steel/Iron Ore 7 0.0510 0.9485 1.9659 0.2068 0.3539 1.4568 0.3633 0.1541 0.8421 0.5944
Hotels/Resorts 5 0.0168 2.2934 3.7945 0.2869 1.1013 4.1555 0.4829 0.8753 3.8558 1.0000
Biotechnology: In Vitro & In Vivo Diagnostic Substances 5 0.0252 0.4688 1.5839 0.2068 0.8971 3.0465 0.4845 0.4424 2.1666 1.0000
Electronic Components 8 0.0234 0.6777 1.6029 0.1435 0.8915 2.7138 0.2949 0.5627 2.2689 0.4826
Average 9.12 0.0291 0.5353 1.2858 0.1658 0.4416 1.7401 0.3297 0.2537 1.2623 0.5932
Table 13. Complete result table for TSPRank with local training on NASDAQ.
Sector Stocks $\tau$ IRR@1 SR@1 MAP@1 IRR@3 SR@3 MAP@3 IRR@5 SR@5 MAP@5
Biotechnology: Laboratory Analytical Instruments 5 0.0403 0.5598 2.6040 0.2236 0.5557 3.3714 0.5270 0.4553 3.1498 1.0000
Diversified Commercial Services 9 0.0411 2.4720 2.5039 0.2954 1.0517 3.0297 0.3104 0.8197 3.1495 0.4093
Other Specialty Stores 11 0.0227 0.0245 0.2376 0.0717 0.6322 1.8056 0.2046 0.4220 1.6928 0.3128
Investment Managers 18 0.0117 1.0477 3.4502 0.0886 0.6783 3.7383 0.1346 0.3774 2.6299 0.1689
Electrical Products 13 0.0218 0.4920 1.4221 0.1181 0.2664 1.3944 0.1587 0.3098 1.7686 0.2540
Major Pharmaceuticals 18 0.0306 -0.1847 -0.0638 0.1814 0.3830 1.4800 0.1917 0.2538 1.4076 0.2024
Other Pharmaceuticals 5 0.0092 0.4138 1.5458 0.2405 0.3732 1.9308 0.4695 0.1719 1.0601 1.0000
Automotive Aftermarket 10 0.0421 0.4458 1.4042 0.1350 0.2815 1.1375 0.2370 0.1293 0.6515 0.3820
Precious Metals 11 0.0319 0.0474 0.3216 0.0844 0.3152 1.3579 0.1988 0.2239 1.1337 0.3240
Publishing 5 0.0277 -0.1257 -0.7956 0.1519 0.0233 0.2224 0.5159 -0.1218 -0.5510 1.0000
Military/Government/Technical 11 0.0371 0.4694 1.4329 0.1266 0.4552 1.9754 0.2239 0.3242 1.8900 0.3363
Business Services 22 0.0333 0.3364 1.1996 0.0464 0.5007 2.9498 0.1036 0.4044 2.9889 0.1477
Packaged Foods 16 0.0406 0.1458 0.5813 0.0928 0.2060 1.0524 0.1685 0.1690 1.0981 0.2349
Life Insurance 17 0.0239 0.0882 0.6505 0.0211 0.2186 1.7630 0.0924 0.2678 1.9882 0.1652
Power Generation 15 0.0353 0.3675 2.5478 0.0591 0.3169 2.3796 0.1531 0.2577 2.3831 0.2139
Clothing/Shoe/Accessory Stores 15 0.0379 2.3676 2.2979 0.1097 0.9053 1.9437 0.1742 0.4445 1.3033 0.2438
Medical Specialities 11 0.0343 0.7813 2.8202 0.1224 0.5672 3.3380 0.2140 0.4546 3.0373 0.3194
Property-Casualty Insurers 30 0.0075 0.0193 0.2590 0.1266 0.1917 0.9142 0.1240 0.1814 1.0815 0.1435
Accident &Health Insurance 5 0.0244 0.3087 1.8361 0.1519 0.3655 2.0095 0.5082 0.2645 1.7726 1.0000
Farming/Seeds/Milling 7 0.0362 0.3846 1.5380 0.1857 0.1224 0.8832 0.3364 0.1262 1.0892 0.5915
Engineering & Construction 5 0.0345 0.1577 0.7986 0.1561 0.2455 1.1889 0.5277 -0.0339 -0.0130 1.0000
Aerospace 9 0.0247 0.7391 1.9837 0.1308 0.5730 2.8210 0.2447 0.4335 2.6282 0.4132
Industrial Specialties 9 0.0156 0.4662 1.8216 0.0928 0.5138 2.5086 0.2639 0.3418 2.0486 0.4138
Specialty Insurers 6 0.0518 0.3274 1.8225 0.2068 0.3344 3.0660 0.3997 0.2600 2.6415 0.7428
Major Chemicals 29 0.0202 0.7541 2.2509 0.0549 0.5005 2.7419 0.0898 0.5291 3.3371 0.1152
Integrated oil Companies 19 0.0167 0.9524 2.2709 0.0886 0.7087 2.8509 0.1526 0.3655 2.0313 0.1893
Air Freight/Delivery Services 8 0.0177 0.4120 1.2900 0.1561 0.4939 2.1393 0.3075 0.2120 1.0822 0.4611
Auto Parts:O.E.M. 18 0.0392 0.2733 0.9433 0.1139 0.1944 0.9760 0.1296 0.3378 1.7314 0.1898
Metal Fabrications 23 0.0090 0.4761 3.1351 0.0169 0.3988 2.1217 0.0830 0.1986 1.1345 0.1143
Professional Services 11 0.0415 0.4349 1.4948 0.1646 0.6978 3.2171 0.2356 0.6156 3.2783 0.3452
Consumer Electronics/Appliances 6 0.0272 0.1768 0.8642 0.2194 0.3083 2.0561 0.4058 0.0786 0.7728 0.7263
Fluid Controls 7 0.0474 0.8668 1.9553 0.2194 0.4234 1.9204 0.3397 0.3069 1.6782 0.6052
Other Consumer Services 17 0.0179 0.5385 1.3441 0.0717 0.2641 1.2278 0.1350 0.5152 2.2771 0.1867
Steel/Iron Ore 9 0.0231 0.7430 1.3593 0.1435 0.5142 1.6196 0.2548 0.3780 1.4982 0.4237
Oil/Gas Transmission 6 0.0322 0.2737 1.7428 0.1519 0.2241 2.0204 0.3973 0.1256 1.2338 0.7301
Plastic Products 6 0.0838 0.5736 2.0377 0.2068 0.5182 3.0642 0.4416 0.4475 3.2359 0.7683
Agricultural Chemicals 7 0.0226 0.1401 1.3763 0.0844 0.0105 0.1532 0.3488 -0.0051 0.0369 0.6279
Containers/Packaging 9 0.0385 0.6422 2.6688 0.2110 0.4324 2.8012 0.2700 0.3217 2.6081 0.4249
Water Supply 7 0.0570 0.3788 1.2039 0.1983 0.3926 2.0605 0.3439 0.4456 2.4154 0.6084
Finance: Consumer Services 20 0.0160 0.9140 2.1926 0.0970 0.4886 2.4532 0.1188 0.1702 1.2730 0.1705
Commercial Banks 21 0.0188 0.0495 0.7665 0.0169 0.1185 1.1677 0.0546 0.1436 1.4201 0.0947
Medical/Dental Instruments 12 0.0101 0.7262 3.7107 0.1013 0.5041 3.1897 0.1854 0.2896 1.9492 0.2843
Computer Software: Prepackaged Software 11 0.0342 0.3481 1.5806 0.0928 0.6486 3.3962 0.1925 0.4644 2.7707 0.3269
Oil Refining/Marketing 9 0.0240 0.8635 2.6404 0.1603 0.1772 1.1876 0.2635 -0.0090 0.0016 0.4166
Hotels/Resorts 10 0.0279 0.8493 3.3624 0.1181 0.3664 2.1728 0.2276 0.3277 2.3952 0.3793
Restaurants 7 0.0654 0.6800 2.9722 0.1435 0.2561 1.9646 0.3556 0.0743 0.6134 0.6355
Department/Specialty Retail Stores 8 0.0213 0.2109 0.7823 0.1561 0.1482 0.7635 0.2872 0.0783 0.5165 0.4944
Marine Transportation 14 0.0413 0.3574 0.9199 0.1519 0.3326 1.4241 0.1847 0.3045 1.6041 0.2481
EDP Services 11 0.0313 0.5444 1.1215 0.2447 0.4464 1.7959 0.2302 0.2479 1.2955 0.3362
Hospital/Nursing Management 10 0.0296 0.4994 1.5208 0.0928 0.4827 1.5102 0.2201 0.3523 1.2736 0.3718
Savings Institutions 6 0.0238 0.0890 0.5546 0.1857 0.2552 1.7050 0.4058 0.1622 1.4018 0.7484
Investment Bankers/Brokers/Service 13 0.0301 0.1306 0.6194 0.1139 0.3248 1.6569 0.1985 0.2423 1.3841 0.2842
Meat/Poultry/Fish 5 0.0244 0.3557 1.7454 0.1350 0.2320 1.7939 0.5115 0.0990 0.8471 1.0000
Beverages (Production/Distribution) 12 0.0224 0.1657 0.7928 0.1055 0.2767 1.8955 0.2189 0.3159 2.6596 0.2953
Homebuilding 16 0.0198 1.2634 2.6371 0.1139 1.1338 4.2467 0.1624 0.8278 3.7611 0.2217
Shoe Manufacturing 5 0.0412 0.6360 1.1819 0.2110 0.5264 1.6604 0.5075 0.3139 1.3117 1.0000
Construction/Ag Equipment/Trucks 6 0.0389 0.5224 1.4423 0.1730 0.5174 2.1491 0.3830 0.5055 2.1878 0.7595
Real Estate 13 0.0156 0.5298 1.8985 0.1097 0.3496 2.1786 0.1953 0.3528 2.5679 0.2669
Package Goods/Cosmetics 10 0.0401 0.5845 1.4541 0.1730 0.4727 2.1772 0.2637 0.2672 1.7438 0.3694
Oilfield Services/Equipment 8 0.0285 -0.1387 -0.4397 0.0928 -0.0712 -0.2511 0.2520 -0.1750 -0.7463 0.5034
Paper 8 0.0297 0.8200 2.5088 0.2194 0.4023 1.8222 0.3010 0.2678 1.4775 0.4899
Mining & Quarrying of Nonmetallic Minerals (No Fuels) 6 0.0490 0.4848 1.1827 0.3080 0.1191 0.6274 0.4039 0.0349 0.2884 0.7267
Railroads 10 0.0199 0.4851 1.4619 0.1350 0.4577 2.4256 0.2414 0.3471 2.0814 0.3683
Home Furnishings 8 0.0198 0.3184 1.0925 0.1899 0.2273 1.0715 0.2979 0.1034 0.6607 0.4743
Apparel 10 0.0177 0.2459 0.9315 0.0970 0.1652 0.8444 0.2199 0.2379 1.2735 0.3689
Office Equipment/Supplies/Services 9 0.0385 0.1088 0.5265 0.1392 0.1758 0.9885 0.2579 -0.0792 -0.3433 0.4275
Forest Products 5 0.0521 0.8712 1.9437 0.2363 0.6435 2.6611 0.5234 0.3948 1.9793 1.0000
Auto Manufacturing 9 0.0460 0.5626 1.7120 0.2068 0.4137 2.4364 0.2571 0.3849 2.3002 0.4122
Services-Misc. Amusement & Recreation 6 0.0585 0.5954 2.6355 0.1772 0.5220 3.4206 0.4334 0.3573 2.7303 0.7524
Semiconductors 5 0.0429 0.6257 2.3595 0.1519 0.5468 2.3792 0.5202 0.3035 1.5702 1.0000
Average 10.97 0.0313 0.5012 1.5710 0.1424 0.3974 1.9735 0.2756 0.2788 1.6662 0.4680
Table 14. Complete result table for TSPRank with local training on NYSE.
Sector Stocks $\tau$ IRR@1 SR@1 MAP@1 IRR@3 SR@3 MAP@3 IRR@5 SR@5 MAP@5
Computer Manufacturing 5 0.0462 0.7131 2.6391 0.1941 0.4412 2.6858 0.5063 0.2167 1.3724 1.0000
Transportation Services 5 0.0790 0.8231 2.6376 0.1772 0.5290 2.1863 0.5211 0.1263 0.7757 1.0000
Medical/Dental Instruments 17 0.0009 0.4891 1.3264 0.0675 0.9051 3.1209 0.1395 0.4674 2.3745 0.1898
Property-Casualty Insurers 19 0.0262 0.5825 1.7370 0.1013 0.1596 0.8014 0.1324 0.1295 0.7147 0.1896
Medical Specialities 5 0.0513 0.3184 1.1598 0.1814 0.0501 0.3507 0.5319 -0.0402 -0.1500 1.0000
Biotechnology: Biological Products (No Diagnostic Substances) 12 0.0339 2.7414 3.5486 0.1435 1.0576 3.2959 0.2161 0.7736 2.9855 0.3079
Telecommunications Equipment 16 0.0361 0.7213 1.6043 0.0928 0.6366 2.5527 0.1320 0.2411 1.3507 0.1929
Medical/Nursing Services 6 0.0406 0.2231 0.8829 0.1857 0.5104 2.0530 0.4067 0.3734 1.6074 0.7533
Water Supply 9 0.0271 0.3681 1.5003 0.0886 0.3183 1.5523 0.2616 0.1904 1.0373 0.4305
Real Estate Investment Trusts 8 0.0606 0.2353 1.3346 0.1266 0.2303 1.6356 0.2996 0.1188 1.0167 0.5047
Business Services 17 0.0327 0.9827 2.0224 0.0886 0.4924 2.0209 0.1296 0.5139 2.5909 0.1958
Major Pharmaceuticals 29 0.0150 1.3801 1.7649 0.0844 1.1300 2.4818 0.0928 0.6112 1.9240 0.1143
Farming/Seeds/Milling 6 0.0557 0.3764 1.2830 0.2152 0.1955 1.0461 0.4126 0.1064 0.7187 0.7526
Industrial Specialties 11 0.0219 0.8028 1.9248 0.1013 1.1003 3.2369 0.2295 0.6864 2.9025 0.3320
Air Freight/Delivery Services 6 0.0182 0.1825 0.6729 0.1646 0.3296 1.3477 0.4015 0.0836 0.4943 0.7340
Electrical Products 13 0.0237 1.3353 2.4725 0.0970 0.8455 2.7356 0.1866 0.5791 2.3875 0.2745
Television Services 8 0.0552 0.8375 2.4448 0.1941 0.3247 1.4238 0.3315 0.0965 0.5797 0.5135
Investment Bankers/Brokers/Service 18 0.0392 1.2698 3.2572 0.0802 0.9476 2.9791 0.1796 0.6891 3.1783 0.2325
Catalog/Specialty Distribution 7 0.0486 2.0857 2.9492 0.1435 0.6505 1.9303 0.3586 0.3549 1.4601 0.6089
Savings Institutions 20 0.0158 0.2971 1.1601 0.0928 0.2722 1.5362 0.1191 0.1138 0.7170 0.1627
Other Consumer Services 9 0.0395 0.7300 1.3770 0.1857 0.2894 1.0383 0.3052 0.2786 1.3046 0.4193
Auto Parts:O.E.M. 6 0.0557 -0.0184 0.0857 0.1857 0.1717 0.8682 0.4147 -0.0410 -0.1008 0.7706
Trucking Freight/Courier Services 11 0.0390 1.0055 1.5900 0.1392 0.6701 1.9763 0.2293 0.5084 1.8487 0.3295
Radio And TV Broadcasting And Communications Equipment 10 0.0496 2.8446 2.7443 0.1603 0.9634 2.7246 0.2543 0.4660 2.0555 0.3792
Biotechnology: Electromedical & Electrotherapeutic Apparatus 5 0.0588 1.7655 3.2478 0.2532 0.7827 2.8964 0.5183 0.3760 1.9405 1.0000
Home Furnishings 8 0.0336 -0.1574 -0.1671 0.1730 -0.0020 0.1388 0.2890 0.1674 0.8236 0.5008
Major Chemicals 7 0.0598 0.7313 2.6136 0.1308 0.2680 1.5711 0.3462 0.2038 1.2628 0.6400
Other Specialty Stores 8 0.0471 0.7342 1.5967 0.1013 0.3793 1.5717 0.2726 0.1491 0.8533 0.5228
Computer Software: Programming, Data Processing 7 0.0482 0.5432 1.8926 0.1814 0.5947 3.3514 0.3317 0.3921 2.4265 0.6468
Restaurants 20 0.0296 0.2455 0.8613 0.0549 0.5447 2.1873 0.1308 0.2523 1.3900 0.1650
Automotive Aftermarket 6 0.0501 0.2125 0.9663 0.1519 0.4733 2.0485 0.4262 0.2650 1.4178 0.7566
Investment Managers 5 0.0664 0.0982 0.5116 0.2152 0.2713 1.6159 0.5342 0.0976 0.7306 1.0000
Oil Refining/Marketing 6 0.0507 -0.1676 -0.5275 0.2869 0.2519 1.3796 0.3964 0.2054 1.2418 0.7240
Real Estate 7 0.0426 0.0887 0.4459 0.1350 0.4726 2.2371 0.3537 0.2984 1.9897 0.5988
Apparel 8 0.0435 0.7246 2.2949 0.1013 0.5383 2.2606 0.2738 0.3009 1.5811 0.5303
Oil & Gas Production 6 0.0754 0.2320 0.6923 0.2321 0.0308 0.2634 0.4196 -0.1448 -0.4294 0.7624
Computer Communications Equipment 5 0.0538 0.6350 1.7178 0.1772 0.1738 0.8243 0.5108 0.0383 0.3025 1.0000
Clothing/Shoe/Accessory Stores 8 0.0516 0.9644 1.4510 0.1646 0.3530 1.0613 0.2832 0.3947 1.2682 0.5028
Commercial Banks 5 0.0656 1.2792 3.0330 0.3080 0.5256 2.3832 0.5127 0.1656 0.9570 1.0000
Hospital/Nursing Management 5 0.0571 1.6740 3.2362 0.3249 0.5750 2.7682 0.5047 0.2216 1.3716 1.0000
Banks 5 0.0681 0.5274 2.1182 0.2321 0.2835 1.5233 0.4913 0.0939 0.6403 1.0000
Professional Services 6 0.0373 0.2523 1.1628 0.1477 0.3277 1.7772 0.4121 0.1111 0.7555 0.7574
Packaged Foods 7 0.0354 0.3166 1.0405 0.1772 0.2745 1.4123 0.3378 0.0144 0.1752 0.5915
Metal Fabrications 12 0.0332 1.3147 2.2192 0.1308 0.5870 2.0254 0.2107 0.4915 2.0341 0.2896
Diversified Commercial Services 6 0.0479 0.8004 1.7045 0.2152 0.6214 2.7428 0.4001 0.3688 2.1574 0.7675
Computer peripheral equipment 6 0.0457 1.1838 2.2897 0.1688 0.9419 3.3506 0.4098 0.3438 1.7482 0.7476
Steel/Iron Ore 7 0.0554 0.5307 1.4018 0.1688 0.4526 1.8570 0.3472 0.2197 1.1517 0.5988
Hotels/Resorts 5 0.0529 1.5454 3.2324 0.2658 1.0232 4.1963 0.5134 0.8753 3.8558 1.0000
Biotechnology: In Vitro & In Vivo Diagnostic Substances 5 0.0588 0.8881 2.3463 0.2152 1.0344 3.4675 0.4930 0.4424 2.1666 1.0000
Electronic Components 8 0.0555 0.9552 1.8544 0.1603 1.1190 3.3023 0.3099 0.6983 2.6961 0.5051
Average 9.12 0.0447 0.7849 1.7471 0.1633 0.5224 2.0359 0.3364 0.2937 1.4331 0.5999
Table 15. Complete result table for TSPRank with global training on NASDAQ.
Sector Stocks $\tau$ IRR@1 SR@1 MAP@1 IRR@3 SR@3 MAP@3 IRR@5 SR@5 MAP@5
Biotechnology: Laboratory Analytical Instruments 5 0.0454 0.5059 2.8110 0.1983 0.4805 3.2350 0.5251 0.4553 3.1498 1.0000
Diversified Commercial Services 9 0.0287 0.6134 1.6824 0.1603 0.7515 2.5416 0.2771 0.5597 2.3811 0.4168
Other Specialty Stores 11 0.0426 0.4883 1.1274 0.1266 0.7789 2.1611 0.2180 0.4915 1.7811 0.3265
Investment Managers 18 0.0319 0.8985 3.0370 0.0844 0.5710 2.9765 0.1282 0.5072 3.0107 0.1797
Electrical Products 13 0.0364 0.5099 1.8447 0.1097 0.3411 1.9502 0.1793 0.3041 1.8406 0.2766
Major Pharmaceuticals 18 0.0484 0.3468 0.8992 0.1392 0.5776 2.5628 0.1620 0.3553 2.0901 0.2052
Other Pharmaceuticals 5 0.0462 0.4623 1.6864 0.2743 0.3309 1.6992 0.4883 0.1719 1.0601 1.0000
Automotive Aftermarket 10 0.0376 0.3244 1.1962 0.1266 0.1897 0.9214 0.2278 0.1262 0.7006 0.3639
Precious Metals 11 0.0423 0.8095 2.1805 0.1688 0.6709 2.6943 0.2344 0.4070 2.0062 0.3355
Publishing 5 0.0504 -0.1658 -0.1896 0.2869 -0.0292 0.0094 0.4887 -0.1218 -0.5510 1.0000
Military/Government/Technical 11 0.0340 0.0927 0.4890 0.1477 0.2599 1.5451 0.2105 0.2256 1.6561 0.3416
Business Services 22 0.0314 0.7742 2.5470 0.0886 0.6654 3.7382 0.1331 0.5076 3.6890 0.1686
Packaged Foods 16 0.0275 0.0566 0.3738 0.0970 0.2547 1.5930 0.1460 0.2264 1.7117 0.2216
Life Insurance 17 0.0325 0.9913 2.8119 0.0886 0.2141 1.4321 0.1486 0.2798 2.0832 0.1960
Power Generation 15 0.0410 -0.0439 -0.0996 0.0675 0.1298 1.1274 0.1535 0.1802 1.5840 0.2119
Clothing/Shoe/Accessory Stores 15 0.0456 1.2433 1.6214 0.1055 0.6930 1.5866 0.1737 0.3455 1.1032 0.2485
Medical Specialities 11 0.0271 0.4398 1.7652 0.0802 0.5340 3.1256 0.1995 0.4774 3.0575 0.3153
Property-Casualty Insurers 30 0.0142 0.2757 0.8266 0.0886 0.2229 1.2077 0.0879 0.1546 1.0534 0.1065
Accident &Health Insurance 5 0.0286 0.1709 1.0124 0.1435 0.2999 1.9586 0.5124 0.2645 1.7726 1.0000
Farming/Seeds/Milling 7 0.0310 0.1738 0.8761 0.1814 0.2130 1.4021 0.3345 0.0954 0.7925 0.5989
Engineering & Construction 5 0.0437 0.2857 1.1477 0.1519 0.3424 1.2229 0.5270 -0.0339 -0.0130 1.0000
Aerospace 9 0.0411 0.6488 2.4841 0.1392 0.4342 2.5592 0.2632 0.3839 2.5218 0.4233
Industrial Specialties 9 0.0348 0.1359 0.6053 0.1561 0.4000 2.5422 0.2689 0.2626 2.1123 0.4247
Specialty Insurers 6 0.0742 0.5958 2.9815 0.2321 0.3912 3.5052 0.4156 0.2610 2.5276 0.7635
Major Chemicals 29 0.0209 0.4391 1.3614 0.0591 0.5311 2.4851 0.0863 0.4246 2.4869 0.1054
Integrated oil Companies 19 0.0298 0.5355 1.4804 0.0759 0.3828 1.8127 0.1322 0.2732 1.6078 0.1797
Air Freight/Delivery Services 8 0.0354 0.8190 2.0955 0.1857 0.4494 2.0238 0.2972 0.2071 1.1508 0.4734
Auto Parts:O.E.M. 18 0.0282 0.6774 2.4992 0.0844 0.2974 1.6579 0.1238 0.2797 1.5649 0.1822
Metal Fabrications 23 0.0315 1.1879 2.4235 0.1266 0.3417 1.4612 0.1414 0.2625 1.3513 0.1663
Professional Services 11 0.0516 1.2433 2.7674 0.1350 0.3915 2.0929 0.2042 0.3667 2.1894 0.3435
Consumer Electronics/Appliances 6 0.0434 0.0869 0.5293 0.1730 0.2392 1.6402 0.3926 0.1074 0.9977 0.7422
Fluid Controls 7 0.0458 0.7086 3.1057 0.1857 0.3542 1.9654 0.3476 0.2994 1.8719 0.6313
Other Consumer Services 17 0.0284 0.2415 0.7103 0.1055 0.0138 0.1909 0.1676 0.3256 1.6478 0.2158
Steel/Iron Ore 9 0.0238 0.9439 1.6691 0.1181 0.3774 1.2905 0.2536 0.3934 1.5161 0.4125
Oil/Gas Transmission 6 0.0580 0.3785 1.7910 0.1772 0.2577 2.1271 0.4144 0.1777 1.6504 0.7406
Plastic Products 6 0.0726 0.3942 1.6841 0.1646 0.5777 3.5004 0.4369 0.3581 2.5870 0.7513
Agricultural Chemicals 7 0.0618 0.0953 0.5161 0.1266 0.2795 1.5774 0.3643 0.0391 0.3648 0.6268
Containers/Packaging 9 0.0362 0.0613 0.5496 0.0759 0.3036 2.4296 0.2196 0.3183 2.6423 0.4258
Water Supply 7 0.0570 0.4097 1.4196 0.2025 0.8891 3.7332 0.3591 0.4585 2.5417 0.6079
Finance: Consumer Services 20 0.0433 0.5376 2.0062 0.0886 0.2829 1.6800 0.1111 0.2882 2.0748 0.1659
Commercial Banks 21 0.0217 0.5504 2.0888 0.0802 0.5011 2.9425 0.1217 0.3375 2.5244 0.1625
Medical/Dental Instruments 12 0.0313 0.1891 1.0436 0.0759 0.4168 2.7722 0.1833 0.4027 2.8406 0.2830
Computer Software: Prepackaged Software 11 0.0261 0.7705 2.1522 0.0844 0.6469 3.1697 0.2011 0.4349 2.3852 0.3323
Oil Refining/Marketing 9 0.0593 0.7053 2.2314 0.1224 0.3295 1.6251 0.2555 0.1965 1.2611 0.4422
Hotels/Resorts 10 0.0402 0.1859 0.9010 0.1181 0.5225 3.0131 0.2403 0.4942 3.4102 0.3747
Restaurants 7 0.0682 0.1222 0.6333 0.1561 0.0546 0.4712 0.3378 0.1528 1.1004 0.6384
Department/Specialty Retail Stores 8 0.0555 0.3472 1.0299 0.1519 0.2370 0.9831 0.2977 0.2523 1.1683 0.5081
Marine Transportation 14 0.0361 0.5846 1.4050 0.1392 0.2815 1.3936 0.1765 0.2895 1.6238 0.2451
EDP Services 11 0.0468 0.6478 1.7467 0.1224 0.2712 1.2282 0.2332 0.2565 1.3250 0.3540
Hospital/Nursing Management 10 0.0399 0.0487 0.3412 0.0464 0.3271 1.2364 0.2157 0.2169 0.9386 0.3695
Savings Institutions 6 0.0546 -0.0320 -0.0924 0.1561 0.2396 1.9658 0.4027 0.1941 1.7806 0.7720
Investment Bankers/Brokers/Service 13 0.0233 0.1814 0.7968 0.1139 0.2372 1.2043 0.1967 0.2408 1.3483 0.2689
Meat/Poultry/Fish 5 0.0723 0.1941 1.0074 0.2068 0.2952 1.8236 0.5122 0.0990 0.8471 1.0000
Beverages (Production/Distribution) 12 0.0435 0.2829 1.4588 0.1266 0.3857 2.8449 0.1906 0.3436 2.9831 0.2874
Homebuilding 16 0.0411 1.2737 2.8299 0.0970 1.0350 3.6869 0.1629 1.0061 4.1459 0.2241
Shoe Manufacturing 5 0.0630 2.4099 2.3840 0.2700 0.6841 1.9317 0.5136 0.3139 1.3117 1.0000
Construction/Ag Equipment/Trucks 6 0.0445 1.1036 2.3375 0.1899 0.5665 2.0795 0.4158 0.3165 1.4641 0.7446
Real Estate 13 0.0218 0.3852 1.8531 0.0464 0.3056 2.0952 0.1261 0.2149 1.7090 0.2573
Package Goods/Cosmetics 10 0.0684 0.5788 2.1845 0.1392 0.4712 2.9352 0.2482 0.3153 2.2773 0.3878
Oilfield Services/Equipment 8 0.0453 0.1043 0.5019 0.1181 0.0309 0.2552 0.2832 -0.0842 -0.2748 0.5072
Paper 8 0.0654 0.1922 0.8538 0.1392 0.3425 1.7754 0.3228 0.2166 1.2878 0.5110
Mining & Quarrying of Nonmetallic Minerals (No Fuels) 6 0.0535 0.0412 0.2960 0.1561 0.2562 1.2813 0.4196 0.0165 0.1869 0.7555
Railroads 10 0.0324 0.0023 0.1227 0.0928 0.3475 2.1375 0.2307 0.1976 1.4698 0.3561
Home Furnishings 8 0.0360 0.5031 1.2453 0.2110 0.3288 1.5796 0.3073 0.2074 1.2672 0.4880
Apparel 10 0.0296 0.0734 0.4790 0.0759 0.3282 1.8800 0.1871 0.3780 2.1074 0.3702
Office Equipment/Supplies/Services 9 0.0479 -0.1323 -0.2453 0.1477 0.1893 0.9815 0.2722 -0.0432 -0.1300 0.4131
Forest Products 5 0.0639 0.9512 2.0830 0.2616 0.6872 2.6944 0.5047 0.3948 1.9793 1.0000
Auto Manufacturing 9 0.0451 0.4649 1.6525 0.1392 0.3408 2.0907 0.2628 0.2689 1.7380 0.4230
Services-Misc. Amusement & Recreation 6 0.0345 0.4000 1.4716 0.2236 0.4410 2.7912 0.4051 0.3083 2.4364 0.7499
Semiconductors 5 0.0571 0.9888 2.7414 0.2068 0.4380 1.9950 0.5087 0.3035 1.5702 1.0000
Average 10.97 0.0422 0.4787 1.4552 0.1392 0.3889 1.9976 0.2756 0.2816 1.7350 0.4732
Table 16. Complete result table for TSPRank with global training on NYSE.
Appendix F MQ2008 Complete Results
Table 17 presents the detailed performance results on the MQ2008 dataset across different folds.
LambdaMART Rankformer TSPRank-Local TSPRank-Global
NDCG@3 NDCG@5 NDCG@10 MRR Tau NDCG@3 NDCG@5 NDCG@10 MRR $\tau$ NDCG@3 NDCG@5 NDCG@10 MRR Tau NDCG@3 NDCG@5 NDCG@10 MRR $\tau$
Fold Top 10
Fold 1 0.6867 0.7180 0.8702 0.4060 0.1330 0.7218 0.7504 0.8832 0.4280 0.2083 0.6837 0.7227 0.8699 0.3956 0.1346 0.7224 0.7606 0.8883 0.4895 0.2174
Fold 2 0.6651 0.7052 0.8631 0.4379 0.1117 0.7057 0.7469 0.8809 0.4784 0.2045 0.6671 0.7031 0.8668 0.4655 0.1193 0.7162 0.7472 0.8838 0.4706 0.1940
Fold 3 0.6904 0.7337 0.8755 0.4286 0.1782 0.7442 0.7743 0.8968 0.4984 0.2651 0.6936 0.7326 0.8735 0.4153 0.1813 0.7341 0.7678 0.8927 0.4883 0.2444
Fold 4 0.6702 0.7180 0.8657 0.4248 0.1406 0.7082 0.7504 0.8814 0.4413 0.2345 0.6930 0.7247 0.8749 0.4662 0.1655 0.7227 0.7519 0.8847 0.4932 0.2192
Fold 5 0.7043 0.7361 0.8791 0.4324 0.1734 0.7301 0.7605 0.8904 0.4844 0.2459 0.6915 0.7232 0.8742 0.3905 0.1714 0.7483 0.7651 0.8926 0.4887 0.2311
Average 0.6833 0.7222 0.8707 0.4259 0.1474 0.7220 0.7565 0.8865 0.4661 0.2317 0.6858 0.7213 0.8719 0.4266 0.1544 0.7287 0.7585 0.8884 0.4861 0.2212
Top 30
Fold 1 0.7284 0.7244 0.7410 0.3563 0.2392 0.7214 0.7221 0.7411 0.3535 0.2497 0.7151 0.7230 0.7401 0.3096 0.2515 0.7587 0.7504 0.7604 0.3685 0.2616
Fold 2 0.7097 0.7088 0.7225 0.3564 0.2061 0.7501 0.7466 0.7598 0.3896 0.2970 0.7189 0.7240 0.7362 0.3206 0.2054 0.7521 0.7534 0.7587 0.3945 0.2725
Fold 3 0.7481 0.7397 0.7463 0.3750 0.2409 0.7641 0.7638 0.7665 0.3860 0.2932 0.7288 0.7304 0.7426 0.3061 0.2444 0.7703 0.7612 0.7637 0.4190 0.2686
Fold 4 0.7386 0.7357 0.7421 0.3754 0.2443 0.7271 0.7292 0.7465 0.3526 0.2591 0.7223 0.7233 0.7368 0.3483 0.2294 0.7547 0.7558 0.7635 0.3833 0.2545
Fold 5 0.7453 0.7406 0.7495 0.3456 0.2554 0.7803 0.7733 0.7841 0.3841 0.3178 0.7122 0.7235 0.7425 0.2997 0.2435 0.7550 0.7583 0.7693 0.3820 0.2675
Average 0.7340 0.7298 0.7403 0.3617 0.2372 0.7486 0.7470 0.7596 0.3732 0.2834 0.7195 0.7248 0.7396 0.3169 0.2348 0.7582 0.7558 0.7631 0.3895 0.2649
Table 17. Complete results on MQ2008-list for information retrieval.
Appendix G OTD2 Complete Results
Table 18 shows the detailed performance results on the OTD2 dataset with five different random group allocations.
LambdaMART Rankformer TSPRank-Local TSPRank-Global
Group Size 10
Allocation $\tau$uparrow EM $\uparrow$ MRR $\uparrow$ RMSE $\downarrow$ $\tau$uparrow EM $\uparrow$ MRR $\uparrow$ RMSE $\downarrow$ $\tau$uparrow EM $\uparrow$ MRR $\uparrow$ RMSE $\downarrow$ $\tau$uparrow EM $\uparrow$ MRR $\uparrow$ RMSE $\downarrow$
1 0.6456 0.3043 0.7813 1.904 0.6269 0.2801 0.7542 1.962 0.5632 0.2909 0.7703 2.369 0.6285 0.3275 0.8038 2.060
2 0.5834 0.2739 0.7504 2.144 0.6196 0.2924 0.7509 2.007 0.5664 0.2682 0.7918 2.226 0.6344 0.3400 0.7997 2.047
3 0.6343 0.3018 0.7534 1.963 0.6144 0.2874 0.7162 2.017 0.5568 0.2917 0.7489 2.356 0.6304 0.3339 0.7751 2.063
4 0.6252 0.3051 0.7431 1.992 0.6123 0.2935 0.7264 2.022 0.5533 0.2632 0.7363 2.274 0.6226 0.3310 0.7903 2.082
5 0.6008 0.2862 0.7538 2.086 0.6217 0.2962 0.7328 1.984 0.5891 0.3138 0.7924 2.255 0.6347 0.3428 0.7993 2.035
Average 0.6297 0.3008 0.7554 1.993 0.6190 0.2899 0.7361 1.998 0.5658 0.2856 0.7679 2.296 0.6301 0.3350 0.7936 2.057
Group Size 30
1 0.5947 0.1066 0.6334 5.948 0.5872 0.0905 0.4825 5.945 0.5026 0.0886 0.5416 6.957 0.6334 0.1378 0.7356 5.729
2 0.5882 0.1013 0.6102 6.047 0.5868 0.0880 0.4927 5.980 0.5256 0.0867 0.6039 6.714 0.6348 0.1351 0.7510 5.728
3 0.5856 0.1135 0.6288 6.071 0.5834 0.0930 0.5059 5.984 0.5116 0.0860 0.5677 6.9300 0.6312 0.1419 0.6895 5.748
4 0.6030 0.1013 0.6062 5.838 0.5863 0.0925 0.4826 5.970 0.5043 0.0873 0.6162 7.0420 0.6290 0.1349 0.7242 5.790
5 0.5930 0.1091 0.5822 5.943 0.5858 0.0967 0.4917 5.986 0.5036 0.0879 0.5401 7.0070 0.6227 0.1423 0.7498 5.857
Average 0.5929 0.1064 0.6122 5.969 0.5859 0.0921 0.4911 5.973 0.5095 0.0873 0.5739 6.930 0.6302 0.1384 0.7300 5.770
Group Size 50
1 0.6074 0.0672 0.5694 9.476 0.5717 0.0543 0.3857 10.084 0.4700 0.0460 0.3961 12.149 0.6179 0.0821 0.6632 9.645
2 0.6072 0.0658 0.5418 9.437 0.5747 0.0523 0.3211 10.012 0.4907 0.0463 0.3514 11.693 0.6189 0.0834 0.6604 9.601
3 0.5973 0.0625 0.5435 9.681 0.5741 0.0523 0.3516 10.040 0.4573 0.0404 0.3529 12.406 0.6169 0.0901 0.6347 9.650
4 0.5989 0.0611 0.5581 9.627 0.5710 0.0541 0.3706 10.086 0.4688 0.0481 0.3617 11.994 0.6231 0.0886 0.6731 9.575
5 0.5891 0.0627 0.5852 9.870 0.5705 0.0505 0.3342 10.122 0.4697 0.0490 0.5122 12.179 0.6268 0.0915 0.6774 9.537
Average 0.6000 0.0639 0.5596 9.618 0.5724 0.0527 0.3526 10.069 0.4713 0.0460 0.3949 12.084 0.6207 0.0871 0.6618 9.602
Table 18. Complete results on OTD2 for historical events ordering.